CASES99: Logbook

CASES99: Chronological Logbook Entries


Return to Logbook Top Page
Entry Date Title Site Author #Graphics
208 Fri 14-Dec-2001Tear-down documentation of radiometersstn1delany
207 Thu 22-Jun-2000T/RH summaryoncley
206 Thu 22-Jun-2000boom angle summaryoncley
205 Thu 22-Jun-2000Prop summaryoncley1
204 Thu 22-Jun-2000Rain summaryoncley
203 Thu 22-Jun-2000Sonic summaryoncley
202 Thu 10-Feb-2000Pyrgeometer configuration in CASES99basedelany
201 Wed 08-Dec-1999Snow cover on pyrgsdelany
200 Fri 03-Dec-1999Roof-top post-CASES99 pyrg-caldelany
199 Wed 01-Dec-1999Pyrgeoemter Campbell coefficientsdelany
198 Tue 30-Nov-1999pyrgeometer coefficientsdelany
197 Fri 19-Nov-1999Cosmos configuration for ops4 & Datel problemsbase
196 Fri 19-Nov-1999ops4 (This project won't end!)basemaclean
195 Fri 19-Nov-1999Baud rates of rad data loggersbasemaclean
194 Thu 18-Nov-1999EVE file downloadsmilitzer
193 Tue 16-Nov-1999final configuration for all ADAMStowerknudson
192 Sat 30-Oct-1999FINIS CASES99delany
191 Sat 30-Oct-1999Daily Status Sat, 30 Oct''99delany
190 Sat 30-Oct-1999End of Operations basejirak
189 Fri 29-Oct-1999moving 10m boomtowerlundquist
188 Fri 29-Oct-1999Daily Statusdelany
187 Thu 28-Oct-1999Licor zerotowerdelany
186 Thu 28-Oct-1999Cleaned radiometer domesdelany
185 Thu 28-Oct-1999New winch cabletowerdelany
184 Thu 28-Oct-1999Daily Statusdelany
183 Wed 27-Oct-1999Sonic problem at Station3stn6delany
182 Wed 27-Oct-1999Licor zerodelany
181 Wed 27-Oct-1999Daily Stautus, 27 Oct''99delany
180 Tue 26-Oct-1999Ragwort ventilationtowerdelany
179 Tue 26-Oct-1999Hauling cabletowerdelany
178 Tue 26-Oct-1999Dirt clodsdelany
177 Tue 26-Oct-1999Cleaned 50 m radiometer domesdelany
176 Tue 26-Oct-1999Ventilation of Ragworttowerdelany
175 Tue 26-Oct-1999Zero Licordelany
174 Tue 26-Oct-1999Cleaning domesdelany
173 Tue 26-Oct-1999Daily Status, 26 Oct''99delany
172 Mon 25-Oct-1999Wind directions on towertowerdelany
171 Mon 25-Oct-1999Daily Status, 25 Oct''99delany
170 Sun 24-Oct-1999Zeroed Licortowerdelany
169 Sun 24-Oct-1999Cleaned radiometer domesdelany
168 Sun 24-Oct-1999Daily Status, 24 Oct''99delany
167 Sun 24-Oct-1999Link to tower down / fiber damagetowersemmer
166 Sat 23-Oct-1999All ADAMS reset at 16:09 localtowersemmer
165 Sat 23-Oct-1999Soil coresdelany
164 Sat 23-Oct-1999Reset Reference Voltagesmiller
163 Sat 23-Oct-1999Cleaning Kryptonsstn3delany
162 Sat 23-Oct-1999Daily Status, Oct 23
161 Fri 22-Oct-1999Cleaning radiometersstn2delany
160 Fri 22-Oct-1999Cleaning Kryptons jirak
159 Fri 22-Oct-1999Replace ATI At Station 6stn6jirak
155 Fri 22-Oct-1999Daily status, Oct 22horst
154 Fri 22-Oct-1999Floppy used spacer used on ragwort lidtowersemmer
153 Thu 21-Oct-1999Comb spacer for Ragwort changedtowerdelany
152 Thu 21-Oct-1999Orientation of sensor boomstowerdelany
151 Thu 21-Oct-1999comb orientation at ragwort changedtowersemmer
150 Thu 21-Oct-1999Daily status, Oct 21horst
149 Thu 21-Oct-1999Daily Maintenancemiller
148 Wed 20-Oct-1999Cleaned 5m KH2O at tower tower
147 Wed 20-Oct-199915m TRH replacedtowersemmer
146 Wed 20-Oct-1999Daily status, Oct 20horst
145 Wed 20-Oct-1999Gsoil at station 1 fixedstn1semmer
144 Tue 19-Oct-1999Began Ops3horst
143 Tue 19-Oct-1999Change krypton gains from 10 to 1 in PAMSsemmer
142 Tue 19-Oct-1999Remounted KH2O sensor on Station 6stn6horst
141 Tue 19-Oct-1999Mounted new h2o/co2 sensor at 20mtowerhorst
140 Tue 19-Oct-1999moved 1.5 sensors to .5mmiller
139 Tue 19-Oct-1999Daily status, Oct 19horst
138 Mon 18-Oct-1999daisy crashed and rebootedtowerhorst
137 Mon 18-Oct-1999Removed 20m h2o/co2 sensortowerhorst
136 Mon 18-Oct-1999Cleaned radiometers at 50m and PAM stationstowerhorst
135 Mon 18-Oct-1999PYG.OUT at station 2 questioned and fixed (?)stn2semmer
134 Mon 18-Oct-1999Daily status, Oct 18horst
133 Mon 18-Oct-1999How to redo covarshorst
132 Mon 18-Oct-1999Rebooted cosmos, ragwort, PAM station ingestbasehorst
131 Mon 18-Oct-1999Covars not runningbasehorst
130 Sun 17-Oct-199920m hot wire installed this afternoon.towerhorst
129 Sun 17-Oct-1999Sonic put back up at pam 6stn6semmer
128 Sun 17-Oct-1999Change ATI at 20 meterssemmer
127 Sun 17-Oct-1999CLOVER crashbasesemmer
126 Sun 17-Oct-1999Krypton back on the air at 20 meterstowersemmer
125 Sun 17-Oct-1999CO2 back on at 20 meters towersemmer
124 Sun 17-Oct-1999bad film data on 16 OCTmiller
123 Sun 17-Oct-1999sonic data at 1.5 and 5 m have interference from hf sheath closuresmiller
122 Sun 17-Oct-1999hfilms down 1600 1630miller
121 Sun 17-Oct-1999Krpyton at 20m off the air at ~ 15:30 local timetowersemmer
120 Sun 17-Oct-1999work on CO2 sensortowersemmer
119 Sun 17-Oct-1999u* from nuw sonics is higher than csat and atik sonics on towerhorst
118 Sun 17-Oct-199920m sonic noisy on v axistowerhorst
117 Sun 17-Oct-1999Daily status, Oct 17horst
116 Sun 17-Oct-1999CD labelsmaclean
115 Sun 17-Oct-1999h2o.k.20m back uptowersemmer
114 Sun 17-Oct-1999h2o.k.20m not workingtowersemmer
113 Sat 16-Oct-1999Tests on cosmos analog channel 123towersemmer
112 Sat 16-Oct-1999Hotfilm sheath signal is backtowersemmer
111 Sat 16-Oct-1999Signals from Hotfilm sheath down for awhiletowersemmer
110 Sat 16-Oct-1999Ragwort back uptowersemmer
109 Sat 16-Oct-1999Status Oct 16horst
108 Sat 16-Oct-1999raising/lowering 40m boomtowerlundquist
107 Sat 16-Oct-1999Cycled power on 10m sonic on 55m towertowerhorst
106 Sat 16-Oct-199920m co2 statustowersemmer
105 Sat 16-Oct-199940m co2 power supply taken downtowersemmer
104 Sat 16-Oct-1999Ragwort off the airtowersemmer
103 Fri 15-Oct-1999moving 40m boomtowerlundquist
102 Fri 15-Oct-1999Daily status, Oct 15horst
101 Fri 15-Oct-1999Cycled power on 10m sonic on 60m towertowerhorst
100 Fri 15-Oct-1999marigold jams networkoncley
99 Thu 14-Oct-1999Moved KH2O from Station 6 to 20m on 60m towerstn6horst
98 Thu 14-Oct-199920m sonic died; replaced w/ATI (and KH2O from Station 6)towerhorst
97 Thu 14-Oct-1999Raising 10m and 40m booms to remove hotwirestowerlundquist
96 Thu 14-Oct-1999Daily status, Oct 14horst
95 Thu 14-Oct-1999No licor cal gasestoweroncley
94 Wed 13-Oct-1999hotwire replacement/look on 55-m tower (Lundquist)towerlemone
93 Wed 13-Oct-1999Daily Statusbasehorst
92 Wed 13-Oct-1999tour up 55m towertoweroncley
91 Tue 12-Oct-1999writing CDsbasemaclean
90 Tue 12-Oct-1999Hotwire archivebasemaclean
89 Tue 12-Oct-1999Replaced hygrothermometer fan at 15mtowerhorst
88 Tue 12-Oct-1999Rpile.out at Station 2stn2horst
87 Tue 12-Oct-1999How to restart adams & evesbasemaclean
86 Tue 12-Oct-1999Tilt of PAM sonics and mean wstn4horst
85 Tue 12-Oct-1999towermiller
84 Tue 12-Oct-1999notes for Steve Semmeroncley
83 Tue 12-Oct-1999Daily Statusbasehorst
82 Tue 12-Oct-1999ati.55m power cycletoweroncley
81 Mon 11-Oct-1999new Ironics card at ragworttowermaclean
80 Mon 11-Oct-1999ragwort overheatingtowermaclean
79 Mon 11-Oct-1999all stations have new krypton gainoncley
78 Mon 11-Oct-199920m H2O/CO2 reinstalledtoweroncley
77 Mon 11-Oct-199920m boom up/CO2 removedtoweroncley
76 Mon 11-Oct-1999Daily Statusbasehorst
75 Mon 11-Oct-1999Cleaned 5m KH2Otowerhorst
74 Sun 10-Oct-1999Nappo pp connected to stns 1&5oncley
73 Sun 10-Oct-199920m ATI was at 10 samp/stoweroncley
72 Sun 10-Oct-1999replaced 10m ATItoweroncley
71 Sun 10-Oct-1999cleaned stn 1-6 kryptonsstn4oncley
70 Sun 10-Oct-199920m boom orientationtowerhorst
69 Sun 10-Oct-1999Daily Statusbaseoncley
68 Sat 09-Oct-1999evening IOP statusoncley
67 Sat 09-Oct-19995m tower down timestowermiller
66 Sat 09-Oct-1999looked at krypton voltagestn1oncley
65 Sat 09-Oct-1999cleaned krypton on stn3stn3oncley
64 Sat 09-Oct-1999T/RH needed new fan stn3oncley
63 Sat 09-Oct-1999maintainance tasksoncley
62 Sat 09-Oct-1999raised and lowered 20m boomtowerlundquist
61 Sat 09-Oct-1999Daily Statusbaseoncley
60 Sat 09-Oct-1999swapped prop at 15mtoweroncley
59 Fri 08-Oct-1999co2.20m fixed itself?toweroncley
58 Fri 08-Oct-1999Moving the 40 m boomtowerlundquist
57 Fri 08-Oct-1999new krypton at 55mtoweroncley
56 Fri 08-Oct-1999Corrections for radiometer deployment errorsstn1horst
55 Fri 08-Oct-1999Daily Statusbaseoncley
54 Fri 08-Oct-1999approximate licor calibrationtoweroncley
53 Fri 08-Oct-1999Rnet cal factorsoncley
52 Fri 08-Oct-1999Licor co2 installedtoweroncley
51 Thu 07-Oct-1999daisy problemstowermaclean
50 Thu 07-Oct-1999ATI sonic gainoncley
49 Wed 06-Oct-1999radiometer workoncley
48 Tue 05-Oct-1999co2 turned offtoweroncley
47 Tue 05-Oct-1999moved rads out of IConcley
46 Mon 04-Oct-1999communication up to stn4stn4oncley
45 Mon 04-Oct-1999fix of T/RH problem?stn2oncley
44 Mon 04-Oct-1999Movement of boomstowerhorst
43 Sun 03-Oct-1999Daily Statusbaseoncley
42 Sat 02-Oct-1999Daily Statusbaseoncley
41 Fri 01-Oct-1999cleaned domesoncley
40 Fri 01-Oct-1999Daily Statusbaseoncley
39 Thu 30-Sep-1999ADAM statustowermaclean
38 Wed 29-Sep-1999ADAM statustowermaclean
37 Wed 29-Sep-1999Krypton spiking problem w/EVE A/Dstn3militzer
36 Wed 29-Sep-1999Tsfc replaced on PAM-3stn3militzer
35 Wed 29-Sep-1999PAM station Configuration Summarybasemilitzer
34 Wed 29-Sep-1999PAM FastOut Port Outages due to EVE-RS485 drivermilitzer
33 Tue 28-Sep-1999playing with rad biases from IConcley
32 Tue 28-Sep-1999ADAM statustowermaclean
31 Mon 27-Sep-1999Daily Statusbaseoncley
30 Mon 27-Sep-1999releveled radiometersoncley
29 Sun 26-Sep-1999ADAM statustowermaclean
28 Sat 25-Sep-1999sonic parameterstoweroncley
27 Sat 25-Sep-1999ADAM status, turned off chem shelter ACtowermaclean
26 Sat 25-Sep-1999Daily Statusbaseoncley
25 Sat 25-Sep-1999Pyg IC started!oncley
24 Sat 25-Sep-1999ADAM statustowermaclean
23 Fri 24-Sep-1999delany
22 Fri 24-Sep-199955m tower guy wire tensionstowerknudson
21 Thu 23-Sep-1999Boom Anglessemmer
20 Tue 21-Sep-1999krypton locations on 60 meter towersemmer
19 Mon 20-Sep-1999PAM Sensor / Station ID Locatormilitzer
18 Sun 19-Sep-1999Changes in CONFIG filessemmer
17 Sun 19-Sep-1999STATUS - Sept 19semmer
16 Sun 19-Sep-1999Tower Prop Pitchssemmer
15 Sat 18-Sep-1999STATUS - Sept 18semmer
14 Wed 15-Sep-1999Installed soil sensorsdelany
13 Tue 14-Sep-1999STATUS - Sept 14semmer
12 Tue 14-Sep-1999Soil sensor installationdelany
11 Mon 13-Sep-1999STATUS: Sept 12semmer
10 Sun 12-Sep-1999STATUS - Sept 11semmer
9 Sat 11-Sep-1999STATUS - Sept 10semmer
8 Fri 03-Sep-1999How to Change PAM Krypton Calibration Filesmilitzer
7 Fri 03-Sep-1999Krypton Calibrations for PAM / Config.dat Filemilitzer
6 Thu 02-Sep-1999Datel sampling of 200 hz nonemaclean
5 Thu 02-Sep-1999new kryptons testedoncley
4 Mon 30-Aug-1999PROP CALSsemmer
3 Thu 29-Jul-1999Sysadmin task listmaclean
2 Fri 18-Jun-1999cosmos quirkymaclean
1 Fri 18-Jun-1999Logbook creatednonemaclean


208: Radiation, Site stn1, Fri 14-Dec-2001 11:35:04 MST, Tear-down documentation of radiometers
This information has been transcribed from my CASES97/99 field notebook.
At tear-down I went around the four surface radiation sites and removed
the radiometers and the soil sensors. I made field notes at the sites.
The notes were made on Saturday, 30 Oct''99, or possible the next day.

Station 1
	  pyrg.in  29260         net 97034     Htfluxplate  87036
	  pyrg.out 26416                       SoilT        001
					       Soilq        001
				       	       Rainguage    13

Station 2
	  pyrg.in  31977         net 94197?    Htfluxplate  943149
	  pyrg.out 31981         K&Zin 940181  SoilT        NCAR 002
	                         K&Zout 940187 Soilq        009
					       Rainguage    3


Station 3
	  pyrg.out  31975        net 94125     Htfluxplate  H983212
	                                       SoilT        NCAR 003
					       Soilq        6
				       	       Rainguage    32

Station 5
	  pyrg.out 31979         net 97035  


207: TRH, Site , Thu 22-Jun-2000 16:47:12 MDT, T/RH summary
General:

Two sensors had problems during CASES99.  Stn3 was noted to have had a bad
fan from 29 Sep-9 Oct (possibly due to the rain on 28 Sep).  15m had a weak
fan and old RH probe, which were replaced on 12 and 20 Oct, respectively.

Stn 5 died due to cable problems which were fixed 4 Oct.  The data that
are present should be good.
 

From logbook:

4 Oct 14:40     Swapped T/RH between stns 2 & 5, Got T/RH working with
                reconfiguration of S-bus cable.

9 Oct 15:59     Replaced T/RH fan on stn3.  Probably bad since at least 29 Sep

12 Oct ~10:00   Replace T/RH fan at 15m, which improved the temperature
                profile.  RH.15m still anomolously low by ~2%.

20 Oct 11:20    Replace entire T/RH at 15m.  "This appears to have increased
                both the T and RH values slightly and they now both fit
                into the 55m tower profile better.  Steve found that the
                previous 50Y was an older type and the PRT had been replaced
                with a smaller, faster-response transducer (during CASES97?)"

206: LOG, Site , Thu 22-Jun-2000 16:32:01 MDT, boom angle summary
This compares the two sets of boom angles which were shot.

TOWER
23 Sep: semmer          21/23 Oct: semmer&delany/jirak  diff (min)
ATIK (10m)      96 44' 35"      97 12 30                28
PROP (15m)      96 58' 10"      97 12 25                14
ATIK (20m)      96 44' 35"      95 48 50                56
PROP (25m)      96 44' 35"      96 21 15                23
CSAT (30m)      96 44' 35"      97 12 20                27
PROP (35m)      96 35' 10"      97 11 55                37
ATIK (40m)      96 44' 35"      96 24 10                20
PROP (45m)      96 44' 35"      96 25 30                19
CSAT (50m)      96 21' 55"      96 12 50                09
ATIK (55m)      96 04' 35"      95 54 20                10

CSAT (0.5m)                     98 24 05
CSAT (2.5m)
CSAT (5m)       96 27' 20"      96 42 30                15


PAM STATIONS
        20 Sep: militzer        26-28 Oct: jirak
stn1    sonic   93 31 20        93 45 43                14
stn1    prop     2               0 48 05                72
stn2    sonic   91 57 57        95 05 23               187!
stn2    prop     0             359 37 40                22
stn3    sonic   88 05 55        88 55 15                49
stn3    prop   358             358 46 35                47
stn4    sonic   90 01 10        89 18 37                43
stn4    prop   357             356 63 00                ??
stn5    sonic   93 03 57        92 32 20                32
stn5    prop     0             358 35 00                85
stn6    sonic   88 26 23        88 06 20                20
stn6    prop   357             356 54 06                 6

205: PropVanes, Site , Thu 22-Jun-2000 15:05:45 MDT, Prop summary
General:

There was one prop failure noted during CASES99 (probably due to rain)
which caused a loss of most data from 15m on the 58m tower between
28 Sep and 9 Oct.  Station 5 lost data due to S-bus cable problems (I think)
which were fixed 4 Oct.

Wind directions from the tower appear not to have been
very accurate.  However, the boom angles were shot twice and differed by only
37 minutes at worst.

Wind directions from sites 1-6 were better behaved, with differences typically
less than 3 degrees.  Site 3 was the worst, with directions consistently about
3 degrees counterclockwise from the average of the rest.  It is possible that
this difference is real (since it was close to a gully).  [Though it was 
site 2's boom angle that was different by 3 degrees during the two shoots -- 
see next entry.]

Sensors:
ASTER props on 58m tower
PAM props on PAMs

Logged problems:
28 Sep 16Z: 15m starts reading 0 much of the time, probably associated with
            a strong rain event -- fixed 9 Oct 16 UTC.
9 Oct 16Z:  15m vane bent back to straight by hand.

Wind direction up the tower:

For most of October, there were constant differences between the props
at different levels of up to 10 degrees, probably caused either by incorrect 
boom angles being entered and/or shot.  The ASTER vanes are easy to bend
(as noted on 9 Oct), but I would expect this to cause a wind-speed dependence
on the direction differences, which isn't obvious to me in the data.

There was a change in this bias on 17 Oct at 45m and 26 Oct at 15m.  The
reason is not known.  The attached plot shows this behavior.






204: , Site , Thu 22-Jun-2000 14:29:50 MDT, Rain summary
General:

From the logbook, there appear to have been 3 significant rain events during 
CASES99:
- Morning of 13 Sep
- Night of 27-28 Sept
- Night of 29-30 Oct.
During the September events, our gauges were not connected to our data system
since their data loggers were in use for the radiometer intercomparison.
The October event wasn't seen in the data, possibly because the storm stopped 
our data systems.  The data from stns 1&2 show small events earlier in October,
but these may be spurious.  Thus, the entire rain gauge data set probably 
is useless.

Details:

We put rain gauges at the inner 3 sites for redundancy (one should have been
sufficient, but we had 3 good ETI gauges).  The inner sites were chosen since
they had radiometer stands, and thus loggers.   No data were taken while 
radiometers were in intercomparison mode (until 5 Oct 14:25), since loggers 
were used for the IC.  A mouse was in the gauge at stn 1 22 Oct.

After that, many light events appear in the data, but total:
stn1: 6.2mm	 stn2: 4.3mm	    
stn3 has glitches in the covars at: 
     99 10 11 181730
     99 10 19 161230
otherwise, it is either NA or 0.

We have only put "rainr" (the rain rate) in the NetCDF files, since "raina"
(accumulation) will depend on the time period requested.

Rain was noted to occur in September, during set-up and the radiometer IC
(when the gauges were not connected), and not again until the very end of 
October.  I.e., the totals seen in the data probably are spurious (perhaps 
bugs caught in the weighing gauges?).  

The logbook entries referencing "rain" are:
13 Sep	    rain all morning, windy all day, clear in the afternoon
28 Sep	    Lots of rain last night with lightning.  Must have had power 
	    outages because the adams all rebooted at the same times.  
	    Toucan reboots cause breaks in the PAM archive.
29 Sep      PAM Stations 5 and 6 had their FASTOUT connection to Toucan/Aster
	    interrupted after a heavy thunder/rain storm 2 nights ago.
	    In both cases the problem was not the cable but the EVE RS232/485
	    converter module located on the EVE electronics interface front 
	    panel.
29 Sep	    After the thunder/rain storm 2 nights ago the old (Everest IRT)
	    sensor was reading 100 deg-C.
1 Oct	    rain predicted to start tomorrow through the weekend
2 Oct	    might have been some rain overnight (but not before we left at 
	    23:30)
16 Oct	    Ragwort was taken off the air last night because of possible
	    rain. The lid has to be left open because of heating. 
16 Oct	    Ragwort is back on the air.  There is a slight chance of rain.
30 Oct	    The end of operations came earlier than we had planned.  A 
	    significant rain event caused the communications to go down. 
	    After talking to Bill Blumen, he said the official end of 
	    operations occured 10/30 at 8:00 a.m.
30 Oct	    Last night there was a severe rainstorm with lots of lightning.

203: Sonics, Site , Thu 22-Jun-2000 13:51:43 MDT, Sonic summary
The following is pieced together from logbook entries and my memory:

When I arrived, one UW didn't remember rotation coefficients.  I reloaded
them.  I think we then put this on stn4.

?We also had problems with an ATI which magically fixed itself when I
worked on it?

9/20:
1	NUW5
2	NUW7
3	NUW3
4	ATI3
5	NUW6
6	ATI2

9/25:
10m:	0505
20m:	0202
40m:	0504
6:      0302
spare:  NUW4?

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	0303?	NUW6	0302	0505	0202	0504	NUW4

9/27: 
- 40m (0504) spiking, removed
- stn4 ATI (3) swapped with spare uw (NUW4?) to have spare ATI (3)

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0302	0505	0202	-	0303
									0504s

~9/29:
- 40m replaced with spare () [must have required 3-way swap with 20m due to hotwire mounts?]

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0302	0505	0303	0202	0504s

10/10: 
- 10m (0505) spiking, replaced with old 40m (0504)
- 20m () was at 10 samp/s, changed to 20 samp/s

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0302	0504s	0303	0202	0505s


10/14: 20m dead (0303), replaced with sonic from stn6 (0302)

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	-	0504s	0302	0202	0505s
									0303d

10/15: 
- Took bad sonic (0303) back to Boulder, 
- Found problem with old 10m (0505) was that it was programmed 
  like a UW array.  Fixed.

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	-	0504s 	0302	0202	0505
									0303[dB]

10/16: blown cap and fuse in 0303? fixed by Gary

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	-	0504s 	0302	0202	0505
									0303[B]

10/17:
- 20m (0302) noisy v, replaced by 0505
- old 20m (0302) with noisy v moved back to stn6

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0302	0504s 	0505	0202	0303[B]


?: 0303 returned to field by Cathy & Tony

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0302	0504s 	0505	0202	0303


10/22: Replaced stn6 (0302) with spare (0303)

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3	NUW4	NUW6	0303	0504s 	0505	0202	0302n


10/27: Stn3 [NUW3?] has bad v&w.  No solution found.  (Probably bad xducer)

1	2	3	4	5	6	10	20	40	spare
NUW5	NUW7	NUW3b	NUW4	NUW6	0303	0504s 	0505	0202	0302n




202: Radiation, Site base, Thu 10-Feb-2000 11:25:28 MST, Pyrgeometer configuration in CASES99
The ten pyrgeoemeters used during CASES99 were set out in two configurations:
 the intercomparison with the pigs mounted side-by side, all up-looking,
 the measurement with the pigs dispersed to their specific location; 
   two up-looking at 50m on the tower boom,
   two down-looking at 50 m on the tower boom,
   two uplooking at 2m, one at station 1, one at station 2,
   four down-looking at 2m, one at station 1, one at station 2, 
                         one at station 3 and one at station 5.
The parameter identifications corresponding to these two configurations are:

           Intercomparison                       Measurement      
            Easterly-most
26416 (Rpile,Tcase,Tdome).out.s02      (Rpile,Tcase,Tdome).out.s2     Bluestem
29260 (Rpile,Tcase,Tdome).in.s02       (Rpile,Tcase,Tdome).in.s2      Bluestem 
31977 (Rpile.Tcase,Tdome).in.s01       (Rpile,Tcase,Tdome).in.s1      Shortgrass
31981 (Rpile,Tcase,Tdome).out.s01      (Rpile,Tcase,Tdome).out.s1     Shortgrass
31975 (Rpile,Tcase,Tdome).out.s05      (Rpile,Tcase,Tdome).out.s5     Weed
31974 (Rpile,Tcase,Tdome).out.a.boom   (Rpile,Tcase,Tdome).out.a.50m  Tower 50m
31980 (Rpile,Tcase,Tdome).in.a.boom    (Rpile,Tcase,Tdome).in.a.50m   Tower 50m
31976 (Rpile,Tcase,Tdome).in.b.boom    (Rpile,Tcase,Tdome).in.b.50m   Tower 50m
31978 (Rpile,Tcase,Tdome).out.b.boom   (Rpile,Tcase,Tdome).out.b.50m  Tower 50m
31979 (Rpile,Tcase,Tdome).out.s03      (Rpile,Tcase,Tdome).out.s3     Drainage
            Westerly-most

The pyrgeometers were in intercomparison configuration:
 1999 September 25, ??:00   to  1999 October 4, ??:00
The pyrgeometers were in measurement configuration :
 1999 October 4, ??:00      to  1999 October 30, ??:00

For more details, including schematics, see:
/net/sssf1/delany/deployments/cases99/lwdiverg/lwdivanal.fm
201: Radiation, Site , Wed 08-Dec-1999 10:34:02 MST, Snow cover on pyrgs
At 1030 Wed 8 Dec''99 I temporarilly depowered the fans to allow a full 
snow cover over all the pyrgs.
It started to snow in the nightime and although the fans had kept the dome 
clear snow had built up on the shield so that the snow formed an annulus 
higher than the top of the domes.
Snow is forecast to continue for the next 24 hours 

The extra snow did not occur, the existing snow on the domes evaporated.
On 9 Dec at 11-30 I restarted the fans and cleaned the domes.

200: Radiation, Site , Fri 03-Dec-1999 15:57:28 MST, Roof-top post-CASES99 pyrg-cal
The boom pigs' data has been plagued by intermittent spikes
Today, at ~1030 I replaced the serial cable from the boom-box Campbel 
to the serial BOB with a new one. This seems to have cured the problem.
At ~1700 I removed all the snow from the radiometers and cleaned the domes.

199: Radiation, Site , Wed 01-Dec-1999 09:15:32 MST, Pyrgeoemter Campbell coefficients
These are the coefficients determined from the oil-bath temp calibration of 
the pyrgeoemters, prior to the CASES99 deployment. These coefficients were 
used during the pre-CASES99 intercomparison in the field and during the 
deployment. They are also to be used for the post-CASES99 intercomparison 
on the roof at the Foothills Lab.
This logbook table was copied from 
/net/sssf1/delany/deployments/cases99/lwdiverg/lbpigcoeff.txt
and then the non-pyrgeoemetrs coefficients were added

     Serial#
        Pile
        CR10 coeff
        Scaled Case coeff B0 x 10^0
        Scaled Case coeff B1 x 10^3
        Scaled Case coeff B2 x 10^6
        Scaled Case coeff B3 x 10^9
        Scaled Dome coeff B0 x 10^0
        Scaled Dome coeff B1 x 10^3
        Scaled Dome coeff B2 x 10^6
        Scaled Dome coeff B3 x 10^9
        Dome/ Case const

                       Connected to the Boom-Box Campbell 
pyg.out.1
    31974
        3.8535
        259.50435
        .00335467
        0.286452
        -2.46374
        3868.8
        .00303966
        .292687
        -7.40681
        2207.35
        3.95
pyg.out.2
    31978
        3.5605
        28031978
        3.5605
        280.85943
        .00335437
        0.287099
        -2.0547
        3531.01
        .00303889
        0.291274
        -5.99237
        1874.43
        4.00
pyg.in.1
    31980
        3.6095
        277.04668
        .00335494
        0.283676
        1.33836
        1980.11
        .00305007
        0.27288
        4.55754
        -137.92
        3.90
pyg.in.2
    31976
        3.852
        259.6054
        .00335494
        0.283467
        1.8651
        1587.36
        .00304992
        0.273924
        3.66477
        18.4609
        3.80



                             Connected to the Sta 5 Campbell
pyg.out
    31975
        4.0275
        248.29299
        .00335599
        .288404
        -6.90295
        6038.71
        .00303124
        0.30936
        -16.7272
        3864.95
        3.05


                               Connected to the Sta 3 Campbell 
pyg.out
    31979
        3.902
        256.27883
        .00335423
        0.287392
        -4.79718
        4588.33
        .00302993
        0.307588
        -15.2242
        3443.34
        3.20


                                Connected to the Sta 1 Campbell
pyg.in
    31977
        3.729
        268.16841
        .00335531
        0.283141
        2.81226
        898.827
        .00304622
        0.280377
        -1.0274
        1105.48
        3.50
pyg.out
    31981
        3.743
        267.16538
        .00335403
        0.283362
        2.6255
        1028.04
        .00304215
        0.286235
        -4.41592
        1743.28
        3.20



                            Connected to the Sta 2 Campbell 
pyg.out
    26416
        (3.570)
        280.11
        .00335452
        0.282295
        2.92866
        954.68
        .00304556
        0.282923
        -3.22195
        1552.84
        (3.56)
pyg.in
    29260
        (3.210)
        311.53
        .00335383
        0.27965
        3.92473
        597.175
        .00306136
        0.253163
        15.4003
        -2315.04
        (3.56)

198: Radiation, Site , Tue 30-Nov-1999 15:45:48 MST, pyrgeometer coefficients
lwdiv
197: ADAMS, Site base, Fri 19-Nov-1999 16:34:34 MST, Cosmos configuration for ops4 & Datel problems


  Cosmos Matrix card s/n 940211101
   16 channel Ironics serial card s/n 129043
   Datel Dvme 601 A/D card s/n 400190
   signal conditioning card # 3
   NCR power supply s/n 000307
   chassis # 4 (missing insulation on enclosure lid)
   Teca cooler # 4 s/n 8940208


At first we had Datel #440690 in this adam. It would work for
about 1/2 hour and then the A/D data (Tbox.cosmos) would quit.
Saw HSTSTA errors on console once.

Datel #440690 was in daisy in the field.
Swapped in Datel #400190 which was in cosmos in the field.  

The adam failed twice over the weekend in this configuration.
(I mxreset it from home once).
On Monday am (10:00 MST) I plugged the adam's power cable into
a different socket on bottom of the power conditioning box.

At 15:30  MST Monday, replaced coax network cable.  The collision
light on cosmos' transceiver was flickering constantly. Jiggling
the coax connector at the adam end would stop the collisions.












196: BaseSystems, Site base, Fri 19-Nov-1999 16:32:36 MST, ops4 (This project won't end!)

ops4 has been created for the post-intercomparison of the radiometers
on the FLAB roof



195: Radiation, Site base, Fri 19-Nov-1999 16:17:42 MST, Baud rates of rad data loggers
NOTE!

The campbell loggers on the PAM stations transmit at 1200 baud.
They were set this way in the field at CASES99 after the 
intercomparison period.
194: PAM-EVE, Site , Thu 18-Nov-1999 11:14:25 MST, EVE file downloads
I uploaded the local flash data from the pams.
It is all concatenated into day packet files:
/net/pam/projects/CASES99/packets/flash/99mmdd.pam
which can be netcdf'd using pdecode (which i have not done).

It appears all files for the days deployed were basically there except
for station 1 which was missing most of them.
A summary of the data is:



CASES99 - EVE Local Storage Summary                                     

Notes                                           
1) Station Data shown in metppf file size in bytes                      
                
2) Station 1 failed to have daily CARDCOPY except on days
   265,266,271-73. Other files were recovered from ramdisk.
   It is assumed that the 'card-lockup' problem likely occurred and
   prevented access to it.  If that happened, then note that the station
   did not get a spurious or human induced reset during the project
   which should have cleared up such a problem.   See the note below
   because the syslog.log file appears to confirm this.
    
                                                
        Station Data                                    
Date            1       2       3       4       5       6
                                                
                                                
metppf.256                                      1088    11696 13-Sep-99
metppf.257              14960                   78336   78336
metppf.258                                      78336   78336
metppf.259              12240   6528    12512   56576   78336 16-Sep-99
metppf.260                      78336   78336   7072    78336
metppf.261                      78336   78336   78336   70176
metppf.262                      74036   71536   78336   78336
metppf.263                      0       78336   78336   78336
metppf.264                      21488   78336   78336   78336
metppf.265      22576           78336   78336   78336   78336
metppf.266      45696   11232   78336   78336   78336   78336 23-Sep-99
metppf.267              55080   78336   78336   78336   78336
metppf.268              272     78336   78336   78336   78336
metppf.269              44290   78336   74364   78336   78336
metppf.270              50264   78336   48672   74278   78336
metppf.271      71646   59328   72182   48672   59122   66787
metppf.272      55414   59122   58188   48672   59328   19604
metppf.273      59328   59328   59328   48672   59328   48672 30-Sep-99
metppf.274              59328   59328   48672   59328   48672 1-Oct-99
metppf.275              59328   59328   48672   59328   48672
metppf.276              59328   59328   48672   59328   48672
metppf.277              59328   59328   46137   59328   48672
metppf.278              59328   59328   48672   59328   48672
metppf.279              59328   59122   48672   59328   48672
metppf.280              59328   59328   48672   59328   48672
metppf.281              59328   59328   48672   59328   48672 8-Oct-99
metppf.282              59328   59328   48672   59328   48672
metppf.283              59328   59328   48672   59328   48672
metppf.284                      59328   48672   59328   48672
metppf.285              57886   59328   48672   59328   48672
metppf.286              59328   59328   48672   59328   48672
metppf.287              59328   59328   48672   59328   48672
metppf.288              59328   59328   48672   59328   48672 15-Oct-99
metppf.289              59328   59328   48672   59328   48672
metppf.290              59328   59328   48672   59328   48672
metppf.291              59328   59328   48672   59328   48672
metppf.292              59328   59122   48672   59328   48672
metppf.293              59328   59328   48672   59328   48672
metppf.294              59328   59328   48672   59328   48672
metppf.295              59328   59328   48672   59328   48672 22-Oct-99
metppf.296              59328   59328   48672   59328   48672
metppf.297              59328   59328   48672   59328   48672
metppf.298      59328   59328   59328   48672   59328   48672
metppf.299      59328   59328   59328   48672   59328   48672
metppf.300      59328   59328   59328   48672   59328   48672
metppf.301      59328   59328   59328   48672   59328   48672
metppf.302      59328   59328   59328   48672   59328   48672 29-Oct-99
metppf.303      59328   59328   59328   48672   59328   48672
metppf.304      40376                                   48672 31-Oct-99
metppf.305                                              48672
metppf.306                                              48672

System Log Files:
                1       2       3       4       5       6
syslog.log      13918   69              1746    4209    24786

1) Station 1 log file is extensive and clearly shows
   "0x14f830 (METPPF-Fil): ERROR: FileToCard: Can't create "/CARD/METPPF.303"
   messages indicating the /card was either:
      - locked up  (most likely) 
      - improperly seated in the pcmcia drive slot
   The flash card itself appeared OK.
   The log file also shows occassional:
       Uninitialized Interrupt!
       Vector number 24 (0-255). Spurious Interrupt
       Program Counter: 0x00003afa
       Status Register: 0x2018
   errors. approximately at (no time tag for these system
       10/15 02:10
       10/20 15:05
       10/24 08:00
       10/28 03:00
       10/30 04:00
       10/30 21:40
   These are hardware errors on the backplane which are probably not the
   /card process itself but my guess is the a/d process, and yes these
   are related to the loss of /card data.

2) Station 2 log file was at the max limit and unfortunately purged on
   the bootup here in Boulder.

3) Station 3 log file, oops, I forgot to copy it and have used that
   flash/box for testing EVE code.

4) Station 4,5 log files were short and rolled over just before the end of
   the project.  Otherwise they showed everything was normal.

5) Station 6 log file was the most extensive.
   It shows 
      10/8 2:52Z, restart
      10/11 22:52, reboot
      10/11/23:45, restart
      10/19, 17:20-20:22, 5 restarts during the period
   Otherwise everything looks good, normal operations.

6) All log files indicate proper GPS clock updates every 3 hours as they
   were supposed to have.



193: ADAMS, Site tower, Tue 16-Nov-1999 11:11:05 MST, final configuration for all ADAMS
final resting place for all ADAM cards at teardown


   Ragwort Matrix card s/n 9006
   16 channel Ironics serial card s/n 821003
   Datel Dvme 601 A/D card s/n 320689
   signal conditioner card # 2
   NCR power supply s/n 000371
   chassis # 1  (missing insulation on enclosure lid)
   Teca cooler # 1 s/n 9121083  NSF 23024


   Daisy Matrix card s/n 9010
   8 channel Ironics serial card (chicory) s/n 014029
   Datel Dvme 601 A/D card s/n 440690
   NCR power supply (belongs to PSU) s/n 002672
   chassis # 4 (missing insulation on enclosure lid)
   Teca coller # 4 s/n 8940208


   Cosmos Matrix card s/n 940211101
   16 channel Ironics serial card s/n 129043
   Datel Dvme 601 A/D card s/n 400190
   Datel Dvme 641 expansion card s/n 001231288
   signal conditioning cards # 3, # 7, and a no number board (belongs to PSU)
   NCR power supply s/n 000307
   chassis - no number (belongs to PSU)
   Teca cooler s/n 9421060 (belongs to PSU)


   Marigold Matrix card s/n 9009
   8 channel Ironics serial card (Edelweiss) s/n 67356
   Datel Dvme 601 A/D card 320589
   NCR power supply (marked dead) s/n 000298
   chassis # 2
   Teca cooler # 3


   Fester Matrix card s/n 9013
   16 channel Ironics serial card s/n 821022
   Datel Dvme 601 A/D card s/n 000940694
   signal conditioning card # 6
   NCR power supply s/n 000238
   chassis # 3
   Teca cooler # 2


192: Status, Site , Sat 30-Oct-1999 16:01:09 CDT, FINIS CASES99
At 1600 the CASES99 data acquisition was completed.

During the afternoon the non-essential sensors were removed from the
tower.
After 1600 ADAMs were powered down and removed
191: Status, Site , Sat 30-Oct-1999 11:31:10 CDT, Daily Status Sat, 30 Oct''99
ISFF/CASES99 Daily status 30 Oct 1999

Weather at 1100 CDT:overcast, 18 degC; light winds  SSE at 5 m/s
Surface: saturated. Last night there was a severe rainstorm with 
lots of lightning.


General: The good weather lasted until just before noon, Fri, 29 Oct.
The frontal system stalled over easter Kansas and the afternoon was
cloudy and very windy.
In the evening a severe rain storm occured and lasted well into
the night. There was lots of lightening and several inches of rain.

The complete data system crashed around 1700 and was brought
up again around 1100 on Sat, 30 Oct. 

It was decided to run the data acquisition to undertake zeroing
and then to terminate data collection.

All the ADAMS have been restarted but Station 5 and 6 resolutely
remain dead.
Cosmos:     all sensors look OK
Ragwort:    all sensors look OK
Daisy:      all sensors look OK
Marigold:   all sensors look OK
Stn1:       all sensors look OK
Stn2:       all sensors look OK
Stn3:       all sensors look OK
Stn4:       all sensors look OK
Stn5:       dead
Stn6:       dead




190: LOG, Site base, Sat 30-Oct-1999 10:48:29 CDT, End of Operations
The end of operations came earlier than we had planned. A significant rain event caused the communications to go down. After talking to Bill Blumen, he said the official end of operations occured 10/30 at 8:00 a.m. We brought up the system in order to archive and shut down the proper way
189: Hotwire, Site tower, Fri 29-Oct-1999 11:33:40 CDT, moving 10m boom
From 10:50-11:10, the 10m boom was raised so that I could remove the hotwire
(#69) at 10m. I had intended to go up to 40m as well, but winds were too
strong.
188: Status, Site , Fri 29-Oct-1999 11:05:50 CDT, Daily Status
ISFF/CASES99 Daily status 29 Oct 1999

Weather at 0900 CDT:clear and windy, 22 degC; strong winds  SSE at 12 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok w&v suspect  ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: Strong winds all night, all systems operating.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.
~

187: CO2, Site tower, Thu 28-Oct-1999 12:30:37 CDT, Licor zero
At ~ 1200 the Licor was zeroed for ~ 15 minutes
186: Radiation, Site , Thu 28-Oct-1999 12:28:53 CDT, Cleaned radiometer domes
Between 1030 and 1100 the radiometer domes were cleaned at Stations 1, 2, 5 and 3. All the fans were operating smoothly
185: Tower, Site tower, Thu 28-Oct-1999 12:26:54 CDT, New winch cable
Between 0930 and 1030 Tony, Cathy, Sean and Julie strung a new winch cable for 
the main tower. 
184: Status, Site , Thu 28-Oct-1999 09:12:31 CDT, Daily Status
ISFF/CASES99 Daily status 28 Oct 1999

Weather at 0900 CDT:scattered high cloud, 12 degC; light winds SW at 2-3 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok    w&v bad   ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: All systems operational except for the v and w of Station 6 sonic anemometer.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
"status" 51 lines, 1980 characters 

183: Sonics, Site stn6, Wed 27-Oct-1999 13:59:30 CDT, Sonic problem at Station3
The UW sonic anemometer v and w were bad for station 3.
At ~1200 all the connectors were checked and the transducers cleaned and 
jiggled slightly.
No change was observed.

182: CO2, Site , Wed 27-Oct-1999 11:21:50 CDT, Licor zero
At 07:25 , 27 Oct''99 Licor zero cannister attached.
At 08:00 , 27 Oct''99 Licor zero cannister removed.
181: Status, Site , Wed 27-Oct-1999 09:27:18 CDT, Daily Stautus, 27 Oct''99
ISFF/CASES99 Daily status 27 Oct 1999

Weather at 0930 CDT: cloudless, 15 degC; light winds S at 3-4 m/s
Surface: dry
IOPs: 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok    w&v bad   ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: Weather perfect. All system fully operational except the sonic
at station 3.
Last IOP has now been completed, but nobody can go home until 08:00 Sunday.
Deserters will be court-martialed and shot.

 
Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

180: ADAMS, Site tower, Tue 26-Oct-1999 20:22:09 CDT, Ragwort ventilation
Inspection at ~1500, 26 Oct''99, indicated that the disc inserted in the morning
to hold open the lid of Ragwort had been placed so that it held the radiation 
shield of the ADAM rather than the lid. This was fortuatous as this spacing 
resulted in good daytime ventilation (max Tbox ~ 32 C for a hot day ). This
arrangement will be maintained for nighttime ventilation
179: Tower, Site tower, Tue 26-Oct-1999 20:15:11 CDT, Hauling cable
Monday, 25 Oct''99, in the afternoon, the thermal imaging equipment was winched
to the 58m level of the tower. During the process the hauling cable snagged in
the winch gears and got chewed up. The wire ran through the boom pulleys and 
fell to the ground wiping out the lowest thermocouple but causing no other 
damage.
On Tuesday Tony and Sean inspected the boom and devised the means to thread a 
new cable.
178: SoilSensors, Site , Tue 26-Oct-1999 20:04:58 CDT, Dirt clods
On 23 Oct''99 dirt clods were taken at the four soil sensor sites. The rough clods were brought to the trailer and were then shaped into prisms,
5 cm deep x ~6cm x ~ 6cm.
The prisms were placed in Al foil dishes and dried atop the oven for three days.
On 26 Oct they were then placed inside the 110C oven to complete their drying 
process.
On 28 Oct the dry weight were measured. 
The clods were waxed and their volumes determined on 29 Oct''99

Station Subsample Dry wt Waxed wt in air Waxed wt in water Volume  Density
  1      A        147.7
         B        135.4
  2      A        179.9
         B        166.0
  3      A        105.0
         B         58.0
  5      A        186.9
         B        179.3
177: Radiation, Site , Tue 26-Oct-1999 20:02:28 CDT, Cleaned 50 m radiometer domes
At ~ 1530 on 26 October the 50 m trolley was pulled in and the four radiometers'
domes were cleaned with water wetted tissue. There was indication of brown dirt
on the domes.
176: ADAMS, Site tower, Tue 26-Oct-1999 13:13:24 CDT, Ventilation of Ragwort
Throughout the nighttime the comb held the lid of Ragwort open ~ 1 cm. 
The temperature only fell to 20C. 
At ~ 0900 the comb was removed and the disc was substituted holding the lid
open ~ 5 cm.
175: CO2, Site , Tue 26-Oct-1999 13:09:08 CDT, Zero Licor
Yesterday between 1515 and 1645 the zero cannister was attached to the Licor
intake. 
Today between 1030 and 1115 the zero cannister was attached to the Licor intake.

The signal during operation was ~  2.36
The signal during zero      was ~ -0.214
174: Radiation, Site , Tue 26-Oct-1999 13:06:08 CDT, Cleaning domes
Between 1030 and 1115 the radiometer domes and the net radiometers were cleaned at Stations 1,3, 5 and 2. Photographs of the ground surfaces under the radiometers were taken.
173: Status, Site , Tue 26-Oct-1999 10:23:29 CDT, Daily Status, 26 Oct''99
ISFF/CASES99 Daily status 26 Oct 1999

Weather at 1000 CDT: clear, 17 degC; light winds SSE at 2 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: All systems up and running. Last night the temperature gradient between the surface
and the 55m level of the tower was ~15 C. There were wind events with complete decoupling
above 25m

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.
~

172: PropVanes, Site tower, Mon 25-Oct-1999 10:18:48 CDT, Wind directions on tower
Inspection of the wind direction plots indicate that the wind direction does not change monotonically with height for the tower.

171: Status, Site , Mon 25-Oct-1999 10:03:20 CDT, Daily Status, 25 Oct''99
ISFF/CASES99 Daily status 25 Oct 1999

Weather at 1000 CDT: Wispy high cirrus, 11 degC; winds SSE at 7 m/s
Surface: dry
IOPs: 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: CASES99 participants enjoyed a good meal at Mascio's Steak and Pasta
         last night.
 
Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

170: CO2, Site tower, Sun 24-Oct-1999 13:01:17 CDT, Zeroed Licor
Between 1230 and 1300 a zero canister was attached to the intake of the 
Licor CO2 sensor.

Front panel display for zero    = - 0.2 
Front panel display for ambient = + 2.4 to 2.5

The pump was running well but I couldn't check the flow as the flow guage
was missing.
Both cal cylinders were empty.
169: Radiation, Site , Sun 24-Oct-1999 12:59:33 CDT, Cleaned radiometer domes
Between 1230 and 1300 the domes of the radiometers at stns, 1,2,5 and 3
were cleaned
168: Status, Site , Sun 24-Oct-1999 11:45:03 CDT, Daily Status, 24 Oct''99
ISFF/CASES99 Daily status 24 Oct 1999

Weather at 1130 CDT: Clear, 10 degC; winds SSE at 4 m/s
Surface: dry
IOPs: IOP #11 scheduled for last night was cancelled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: 
  - At 0315, 24 Oct''99 one strand of the main fiber optic connection between 
  the network hub at the base of the tower and the data acquisition system 
  in the trailer was severed. The connection was reestablished at 0930. 
  No data was acquired during the period, 0315 - 0930.
  Rodent involvement is suspected rather than that of a rival gang 
  of micrometeorologists. 

Notes:
"up" - Dep       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 
clover:$ASTER/projects/CASES99/doc: 
loyed and transmitting data, b

167: Network, Site tower, Sun 24-Oct-1999 09:45:49 CDT, Link to tower down / fiber damage
The link to the tower went down at 3:15 am local.
The problem was a bad fiber in the 300m cable. The
bad fiber is orange. All stations were back on the
air at 9:50 am local.
The 300m cable was checked for cuts. Nicks were
found at 2 locations.
We are now down to 4 fibers with 1 week to go.

166: ADAMS, Site tower, Sat 23-Oct-1999 17:04:27 CDT, All ADAMS reset at 16:09 local
All the ADAMS were reset at 16:09 local.
At this time, I do not know why. They came back
up without the need for human input.


165: SoilSensors, Site , Sat 23-Oct-1999 14:27:31 CDT, Soil cores
Core samples were taken at stations 1, 2, 3 and 5, in the vicinity of
the soil sensors. The 0-5 cm portion of the cores were separated, mixed and 
sampled in triplicate. The subsamples were weighed fresh and then after 24 hr
in the drying oven at 110 C.

Corer (TD) Fresh taken 1300, 23 Oct, dry weight measured on 26 Oct.
date/time(CDT) Stn Subsample Fresh  Dried Diff Grav fract
23 Oct 1300     1    I        21.0  18.3
                     II       23.2  20.2
                     III      19.4  16.9
                2    I        20.0  18.5
                     II       21.2  19.6
                     III      21.0  19.3
                3    I        27.0  16.7
                     II       19.4  11.2
                     III      21.0  12.5
                5    I        21.6  20.2
                     II       21.7  20.4
                     III      26.6  24.9

Corer (TD), Fresh taken 1130, 29 Oct, dry wt measured .........
date/time(CDT)  Stn Subsample Fresh   Dried Diff Grav fracr
29 Oct 1130      1   I        19.9     17.6
                     II       22.9     20.5
                     III      20.2     18.0
                 2   I        16.6     15.4
                     II       17.2     15.8
                     III      18.5     17.0
                 3   I        17.4      9.8
                     II       17.2      9.9
                     III      17.8     10.3 
                 5   I        23.5     21.8
                     II       21.5     19.9
                     III      22.7     21.1


Where Grav fract = gravimetric fraction = (fresh wt - dry wt) / (dry wt)

P.S. (by oncley)

Processing these data yields:
Day		Stn				Average Gravimetric
23 Oct		1				0.148
		2				0.084
		3				0.676
		5				0.067
29 Oct		1				0.123
		2				0.085
		3				0.747
		5				0.078
(still need bulk density to compare to volumetric measurements)

164: Hotfilm, Site , Sat 23-Oct-1999 12:25:22 CDT, Reset Reference Voltages
23 Oct
1200 - 1205 reset bridge reference voltages to 5.00 Volts
1206 performed zero wind calibration for the remainder on the hour using the reset reference voltages.
1207 running


163: Kryptons, Site stn3, Sat 23-Oct-1999 11:53:45 CDT, Cleaning Kryptons
Between 1100 and 1145, the Krypton hygrometers at stations 2, 5, 4 and 3 
were cleaned using water wetted Q-tips and were then dried using Q-tips
162: Status, Site , Sat 23-Oct-1999 10:07:20 CDT, Daily Status, Oct 23
ISFF/CASES99 Daily status 23 Oct 1999

Weather at 1000 CDT: Clear, 8 degC; winds NNE at 8 m/s
Surface: dry
IOPs: IOP #11 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General:
  - Bad sonic at station 6 has been replaced. 

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.



161: Radiation, Site stn2, Fri 22-Oct-1999 13:33:22 CDT, Cleaning radiometers
At  11:15 the domes of the radiometers at station 1 were cleaned
At ~13:00 the domes of the radiometers at stations 2, 3, and 5 were cleaned.

There were traces of brownish grey dust, especially on station 1 domes

All net radiometers were cleaned.
All fans were running. The fan at station 3 was noisy.
160: Kryptons, Site , Fri 22-Oct-1999 11:51:50 CDT, Cleaning Kryptons
Cleaned Kryptons 1 and 6. Radiometer domes on station 1 were cleaned. Significant road dust was on the domes. Mouse also in the rain gauge.
159: Sonics, Site stn6, Fri 22-Oct-1999 11:47:23 CDT, Replace ATI At Station 6
Noted yesterday the V component of the ATI at station 6 had a positive offset compared to rest of the sonics. Presious status has noted peculiarities in the response of this sonic. Replaced with the sensor brought out by TOny and CAthy on the 20th. Old sensor # is ATI 980302 (ATI#2). New Sensor ATI 98303.

155: Status, Site , Fri 22-Oct-1999 10:36:41 CDT, Daily status, Oct 22
ISFF/CASES99 Daily status 22 Oct 1999

Weather at 1030 CDT: Clear, 16 degC; winds NE at 10 m/s
Surface: dry
IOPs: IOP #10 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	BAD	ok      -       -       ok      ok      ok 	-

General: 
  - No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


154: ADAMS, Site tower, Fri 22-Oct-1999 10:15:22 CDT, Floppy used spacer used on ragwort lid
The ragwort lid spacer, the comb, was repalced with a
floppy disk at ~10:00 local.


153: ADAMS, Site tower, Thu 21-Oct-1999 17:40:46 CDT, Comb spacer for Ragwort changed
At 1730 the comb spacer for Ragwort was changed from vertical to horizontal.
Note that even when vertical the spacing was not large because plastic comb
was deformed until nearly bent double.
152: Tower, Site tower, Thu 21-Oct-1999 16:36:22 CDT, Orientation of sensor booms
Steve Semmer and Tony Delany shot the angles of the ten east-side booms
between 1500 and 1630 CD Time, Thurs, 21 Oct. The tripod was set up on
the road. The 5m tower sonic anemometers could not be accurately viewed
from that distance.

    55m  Sonic  275 54' 20"     
    50m  Sonic  276 12' 50"
    45m  Prop   276 25' 30"
    40m  Sonic  276 24' 10"
    35m  Prop   277 11' 55"
    30m  Sonic  277 12' 20"
    25m  Prop   276 21' 15"
    20m  Sonic  275 48' 50"
    15m  Prop   277 12' 25"
    10m  Sonic  277 12' 30"

10/23 Cathy and Steve shot the angles of the 2 sonics on the 5 metre tower.
The bottom sonic (.5m) seemed to be unlevel as well as at an angle to the boom.

     5m CSAT	276 42' 30"
    .5m CSAT    278 24' 05"


151: ADAMS, Site tower, Thu 21-Oct-1999 11:53:21 CDT, comb orientation at ragwort changed
The high today is predicted to be high 70s. So Cathy
and I changed the comb orientation from horizontal
to vertical.

150: Status, Site , Thu 21-Oct-1999 11:52:14 CDT, Daily status, Oct 21
ISFF/CASES99 Daily status 21 Oct 1999

Weather at 1015 CDT: Clear, 15.5 degC; winds WSW at 2 m/s
Surface: dry
IOPs: IOP #9 finished this morning; 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	ok	ok      -       -       ok      ok      ok 	-

General: 
  - Fixed Gsoil at Station 1

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Sensor data obviously bad
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


149: Hotfilm, Site , Thu 21-Oct-1999 11:25:19 CDT, Daily Maintenance
21 Oct
1110-1120 CST Reset all reference voltages to 5.00 Volts


148: Kryptons, Site tower, Wed 20-Oct-1999 14:56:16 CDT, Cleaned 5m KH2O at tower
I cleaned the windows of the KH2O at 5m at the main tower, using
a q-tip and distilled water.  I intended to also clean those on the
PAM stations, but realized that the short ladder was not adequate
(duh!).  Since I was too lazy to shlep the large ladder around to
all the PAM stations, I bagged the whole thing.

147: TRH, Site tower, Wed 20-Oct-1999 11:37:37 CDT, 15m TRH replaced
The 15m TRH humidity level has been low compared to the
other sensors on the tower, ~ 2%. The sensor was replaced
with one of the spares. This occurred at ~11:20 local time.
Will monitor its performance.

1500 CDT, Oct 20
This appears to have increased both the T and RH values slightly
and they now both fit into the 55m tower profile better.  Steve
found that the previous 50Y was an older type and the PRT had
been replaced with a smaller, faster-response transducer (during
CASES97?) -- twh


146: Status, Site , Wed 20-Oct-1999 10:42:42 CDT, Daily status, Oct 20
ISFF/CASES99 Daily status 20 Oct 1999

Weather at 1030 CDT: Clear, 12 degC; winds SW at 8 m/s
Surface: dry
IOPs: IOP #8 finished this morning; IOP#9 tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	ok	ok      -       -       ok      ok      ok 	-

General: 
  - Changed PAM KH2O gain back to 1
  - Replaced ANL H2O/CO2 sensor at 20m
  - Moved KH2O back to Station 6
  - Fixed Gsoil at Station 1

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Sensor data obviously bad
X    - Sensor is missing.  
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


145: SoilSensors, Site stn1, Wed 20-Oct-1999 10:21:27 CDT, Gsoil at station 1 fixed
The Gsoil cable was providing food for some hungry creatures.
The bad section of cable was removed and a splice was made.
The sensor looks good now.
NOTE: The change in cable may affect the calibration. data should
be viewed with caution.

144: BaseSystems, Site , Tue 19-Oct-1999 19:43:03 CDT, Began Ops3
Stopped all systems to begin ops3.  Ops3 was needed to change the
gain on the PAM A/D.

PAM stations at 0022 GMT, Oct 20
cosmos       at 0037 GMT
daisy        at 0040 GMT
marigold     at 0041 GMT
ragwort      at 0042 GMT
143: Kryptons, Site , Tue 19-Oct-1999 19:23:37 CDT, Change krypton gains from 10 to 1 in PAMS
Tom decided it was better to get some data with poor
resolution than no data at all! So we changed the
gain setting for the kryptons back to 1.
  A new set of krypton.dat files were loaded into
each PAM station so if we want to change back to
a gain of 10 the files are ready to go.
file names:  Kxxxxg1.dat  - gain of 1
             Kxxxxg10.dat - gain of 10
where "xxxx" is the sensor serial number.
NOTE: only the files for the sensor at the PAM
station have been downloaded.

This change was made from 1530 to 1615 CDT, Oct 19.

142: Kryptons, Site stn6, Tue 19-Oct-1999 19:11:56 CDT, Remounted KH2O sensor on Station 6
Replaced the KH2O sensor on Station 6.  This had been mounted at 20m
on the 55m tower for several days.  We lowered the mast around 1835 CDT
and raised it again around 1900 CDT, Oct 19.

This means that the h2o.k.20m on the tower no longer exist!
However, it is still in the config files for ops3.


141: CO2, Site tower, Tue 19-Oct-1999 19:07:45 CDT, Mounted new h2o/co2 sensor at 20m
ABLE (Argonne) provided a replacement for the bad h2o/co2 sensor
at 20m.  We mounted this on the tower this evening, raising the 20m
boom around 1710 CDT and lowering it at 1755 CDT.

At this time we removed the KH2O sensor for remounting on Station 6.



140: Hotfilm, Site , Tue 19-Oct-1999 14:18:36 CDT, moved 1.5 sensors to .5m
19 Oct
12:51 tried to blow the cobwebs off the cold film and triple film at 1.5m - partially successful

1300-1400  lowered 1.5 m arm to 0.5m.  Releveled sensors - sonic wouldn't level completely - leaning slightly sideways - toward true north. It appears that the weight of the hf holders on that side has bent the arm slightly.

had to remove the tube from the Spanish pressure sensor to lower the arm.  Resealed the connection with silicone at 13:55.

13:44 Reset all reference voltages on the bridges to 5.00 v. 


139: Status, Site , Tue 19-Oct-1999 11:21:28 CDT, Daily status, Oct 19
ISFF/CASES99 Daily status 19 Oct 1999

Weather at 1100 CDT: Mostly clear, 12 degC; winds NW at 5 m/s
Surface: dry
IOPs: IOP #8 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD     ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     ok      -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	ok	X       -       -       ok      ok      ok 	-

General: No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - The 20m h20/co2 is still not working.  ABLE will provide a 
       replacement from their Smileyburg site.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  We will change the range today.  
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing.  We will replace the KH2O at Station 6.
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


138: ADAMS, Site tower, Mon 18-Oct-1999 18:01:18 CDT, daisy crashed and rebooted
Daisy stopped around 16:10 CDT, Oct 18.  We could not 'ping' daisy,
so Steve recycled power at 17:59 CDT.  It was around 16:10 that
we were working on the tower, but it is not apparent how we could
have caused this crash.  We lowered the 20m boom at 16:05 and then
went up to 50m to move the radiometers to the end of their boom.
137: CO2, Site tower, Mon 18-Oct-1999 16:40:30 CDT, Removed 20m h2o/co2 sensor
We have given up on fixing the ANL h2o/co2 sensor on the 20m boom.
We raised the boom at 15:20 to remove it and lowered the boom at
16:05 (after adding another collar to hold the boom pivot in 
position).
136: Radiation, Site tower, Mon 18-Oct-1999 16:38:11 CDT, Cleaned radiometers at 50m and PAM stations
We retracted the 50m radiometers and cleaned them, beginning at 1500 CDT,
Oct 18.  They were cleaned with "de-mineralized" water and kimwipes.

The radiometers were returned to their outboard position at 1615 CDT
(after unfouling the hauling line with the loooon green hook).

The domes of the surface radiation sensors were also cleaned after 
finishing the tower radiometers.


135: Radiation, Site stn2, Mon 18-Oct-1999 11:59:54 CDT, PYG.OUT at station 2 questioned and fixed (?)
The pyg.out pile was bad so I went out to look at it.
According to Tom, this problem has happened before.
The voltage going to the multiplexer board was varying
quite a bit. I played with the connector inside the Campbell
which resulted in a good signal. The cables had been reversed
from a previous test so I switched them back to the correct
configuration. Will need to monitor to see if fix works. If
so, we will need to flag this Campbell box for repairs.

134: Status, Site , Mon 18-Oct-1999 11:36:19 CDT, Daily status, Oct 18
ISFF/CASES99 Daily status 18 Oct 1999

Weather at 1130 CDT: Mostly clear, 12 degC; winds SW at 3 m/s
Surface: dry
IOPs: IOP #7 ran from 1700 Oct 17 to 0800 Oct 18

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      X       ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     Rlw.o   -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	ok	X       -       -       ok      ok      ok 	-

General:
 - The 20m h20/co2 is still not working.  We will return it to ANL.
 - We have put a new sonic at 20m and returned the sonic to station 6  

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy.  We are replacing the power
       supply and are optimistic that this will fix it.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  If we change the range, the resolution is poor.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
Rlw.o- The outgoing long wave radiometer at station 2 is missing Rpile.
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 




133: BaseSystems, Site , Mon 18-Oct-1999 11:03:30 CDT, How to redo covars
In order to redo recently lost covars, edit
$ASTER/projects/CASES99/scrips/covar.redo with an appropriate start time.
Then execute covar.redo.
132: ADAMS, Site base, Mon 18-Oct-1999 10:19:11 CDT, Rebooted cosmos, ragwort, PAM station ingest
Rebooted cosmos and ragwort at 15:18 GMT in order to properly restart
covars.

Also stopped and restartd serial ingest of PAM data on toucan at 
15:20 GMT for the same reason.
131: BaseSystems, Site base, Mon 18-Oct-1999 09:00:07 CDT, Covars not running
The covars quit around 17:25 CDT Oct 17.  Those for daisy and marigold
were restarted following the hangup with Clover, but the others were not 
restarted until 0900 CDT oct 18.
130: Hotwire, Site tower, Sun 17-Oct-1999 19:39:08 CDT, 20m hot wire installed this afternoon.
17 Oct
The 20m boom was raised this afternoon, Oct 17, at 13:13:30 CDT.  
It was lowered at 13:23 CDT.  At this time the hot wire was installed.



129: Sonics, Site stn6, Sun 17-Oct-1999 19:28:35 CDT, Sonic put back up at pam 6
Tom installed the sonic back on station 6.  The mast was back in its
upright position at 17:15 CDT, Oct 17.



128: Sonics, Site , Sun 17-Oct-1999 19:23:54 CDT, Change ATI at 20 meters
Tom and I changed sonics at the 20 meter height. The boom was lowered
at 18:25 CDT, Oct 17.
The sonic at 20 meters is 980505. The sonic that came down went back
to station 6 where it came from a few days ago.



127: BaseSystems, Site base, Sun 17-Oct-1999 19:20:51 CDT, CLOVER crash
While Tom and I were installing the CO2 sensor at 20 meters,
clover suffered a crash. Dave Miller was taking credit for it!
I had to power cycle clover to get it going again. This crash
also shutdown the archiving of blazing data. After a short conversation
with Gordon, marigold and daisy were mxreset to get things going again.

126: Kryptons, Site tower, Sun 17-Oct-1999 19:19:37 CDT, Krypton back on the air at 20 meters
The krypton at 20 meters is back on the air. It is the
primary h2o sensor. The CO2 was also remounted and is
on the air. The boom was lowered at 18:25 CDT.




125: CO2, Site tower, Sun 17-Oct-1999 19:17:25 CDT, CO2 back on at 20 meters
The CO2 sensor was remounted at 20 meters after a preliminary
test with the new 5 volt supply. We left the krypton there
as the primary h2o sensor. The boom was lowered at 18:25 CDT.



124: Hotfilm, Site , Sun 17-Oct-1999 18:11:12 CDT, bad film data on 16 OCT
16 Oct
Checked data for previous day.  Bad film (vu.hf.5m)  on 16 Oct from about 1000 to 2200 local time.



123: Sonics, Site , Sun 17-Oct-1999 16:46:37 CDT, sonic data at 1.5 and 5 m have interference from hf sheath closures
The hot film sheath closing every hour for 1 minute can be seen in the variances of the adjacent sonics at 1.5 and 5 m. So the sonic data during the sheath closures is suspect.



122: Hotfilm, Site , Sun 17-Oct-1999 16:41:37 CDT, hfilms down 1600 1630
17 Oct
1600 - 1630 reset hf/cf bridge voltages and checked levels.  All sensors are leaning slightly to the north (ie level bubble still in central circle).




121: Kryptons, Site tower, Sun 17-Oct-1999 14:49:42 CDT, Krpyton at 20m off the air at ~ 15:30 local time

120: CO2, Site tower, Sun 17-Oct-1999 14:46:38 CDT, work on CO2 sensor
It appears that the problem with the CO2/H2O sensor was
in the 5 volts supply. I modified their power supply so
that we could use a different 5 volt source. The sensor
was deployed for testing at the 20 meter height. After
we collect some data we will decide if it is good or bad.
The sensor was deployed at ~13:50 local time.

119: Sonics, Site , Sun 17-Oct-1999 13:32:22 CDT, u* from nuw sonics is higher than csat and atik sonics on tower
With the high winds, u* from the nuw sonics is noticeably higher than
u* for all the sonics on the 55m tower, including those at 1.5, 5 and 
10m.
118: Sonics, Site tower, Sun 17-Oct-1999 13:03:14 CDT, 20m sonic noisy on v axis
The v axis on the 20m ATIK has been spiking.  We plan to move this unit
back to station 6 this afternoon and mount the sonic removed from 
10 m on Oct 10.  SPO found that this sonic was improperly programmed
as though it were a UW array.  He fixed this and we will mount it at
20m.

117: Status, Site , Sun 17-Oct-1999 12:30:10 CDT, Daily status, Oct 17
ISFF/CASES99 Daily status 17 Oct 1999

Weather: Clear, 12 degC; winds NE at 8 m/s;
Surface: dry
IOPs: IOP #7 scheduled to begin at 1700

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      X       -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  v?      ?       X       X       -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       X       -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     ok      -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - Ragwort restarted at 14:40 CDT, Oct 16
 - We hope to get the 20m h20/co2 working today.  If so, we will return the
   KH2O to station 6.  We also have a replacement sonic for station 6.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy.  We are replacing the power
       supply and are optimistic that this will fix it.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  If we change the range, the resolution is poor.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 



116: BaseSystems, Site , Sun 17-Oct-1999 10:26:58 CDT, CD labels

Tom,

I Fedex'd 100 CD labels this morning.  The Boulder store also did
not have the identical product.  I bought some Memorex
labels that are perforated in the middle and were folded in
half in their box. Otherwise they looked the same as the others.
If you feed them in whole they should work.

If you run out of labels before they arrive, don't use a pen
to label the CD's.  You're supposed to use a special CD pen -
otherwise some inks can apparently leach through the plastic
and spoil the substrate.

If you forget what is on a CD, you can just insert it into
one of the CD drives on clover or cocklebur (except the CD-R drive).
A file manager window should popup, which will allow you
to browse the CD, and then eject it when you're done.

If a window doesn't popup, you can use shell comands to look
in the /cdrom directory.  Afterwards, cd out of /cdrom
and do:
	eject cdrom


I notice that marigold got in a state where it flooded the network
with packets.  This happened once in early Oct too.  I don't
know the reason or any fix, except rebooting.

Gordon

115: Kryptons, Site tower, Sun 17-Oct-1999 10:01:46 CDT, h2o.k.20m back up
The krypton at 20m is back on the air. The analog
signal cable was plugged into the wrong channel.

114: Kryptons, Site tower, Sun 17-Oct-1999 09:35:09 CDT, h2o.k.20m not working
The 20m krypton was not working throughout the night.
I may have incorrectly connected it back up after the
breakout box test.

113: ADAMS, Site tower, Sat 16-Oct-1999 16:13:04 CDT, Tests on cosmos analog channel 123
During the day, tests were run to determine the source of noise
on analog channel 123, h2o.k.20m, on cosmos. Below is the list of
tests and the outcome.

(1) switched h2o.5m with h2o.k.20m. So h2o.5m signal was going
into channel 123 and h2o.k.20m was going into 108.
result: noise problem stayed with channel 123.

(2) changed analog/filter board
result: noise on 123

(3) removed special control board near power supply
result: noise on 123

(4) reversed h2o.5m and h2o.k.20m. switched analog breakout boxes
result: noise still on channel 123

112: Hotfilm, Site tower, Sat 16-Oct-1999 16:12:28 CDT, Hotfilm sheath signal is back

111: Hotfilm, Site tower, Sat 16-Oct-1999 15:02:18 CDT, Signals from Hotfilm sheath down for awhile
16 Oct
The signals from the hotfilm sheath are down. They will be
down for a test of an analog breakout box.



110: ADAMS, Site tower, Sat 16-Oct-1999 15:01:25 CDT, Ragwort back up
Ragwort is back on the air. Tom and I decided to leave it
on tonight. There is a slight chance of rain.

109: Status, Site , Sat 16-Oct-1999 11:56:48 CDT, Status Oct 16
ISFF/CASES99 Daily status 16 Oct 1999

Weather: Cloudy, cooler; winds N at 9 m/s;
	 frontal passage around 2230 CDT, Oct 15.
Surface: dry
IOPs: None scheduled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      X       -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ?       X       X       -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       X       -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        shut down at 1730 CDT, Oct 15, see (r) above
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - Ragwort has a bad cooler.  We have left the lid open to compensate,
   but shut the lid last night in anticipation of rain and turned off
   ragwort to avoid internal damage due to heating.  Will replace 
   cooler Sunday or Monday.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 




108: Hotwire, Site tower, Sat 16-Oct-1999 09:25:32 CDT, raising/lowering 40m boom
At 08:47 CDT, I raised the 40m boom to uninstall the #70 hot wire from the
intercomparison experiment last night with HRDL. By 08:53, it was back 
down. 
107: Sonics, Site tower, Sat 16-Oct-1999 09:20:04 CDT, Cycled power on 10m sonic on 55m tower
We turned off cosmos temporarily at 1350 CDT, Oct 15, for a test of
the filtering on the 20m KH2O.  This stopped the 10m sonic again.
Power was cycled on the 10m sonic at 1715 CDT to restart it.
106: CO2, Site tower, Sat 16-Oct-1999 08:46:59 CDT, 20m co2 status
I ran some tests on the 20m co2 sensor. Based on feedback from
Dave Cook at Argone, it appears the sensor is bad. At one time,
it appeared the 5 volt supply was the problem; however, tests
with the other co2 supply showed this not to be the case. We will
need to send the unit back to Dave Cook.

105: CO2, Site tower, Sat 16-Oct-1999 08:43:48 CDT, 40m co2 power supply taken down
The power supply unit for the 40m co2 sensor was removed
last night to run tests with the questionable co2 sensor
deployed at 20m. No data is being lost since ragwort has been
turned off.


104: ADAMS, Site tower, Sat 16-Oct-1999 08:41:09 CDT, Ragwort off the air
Ragwort was taken off the air last night because of possible
rain. The lid has to be left open because of heating. Hopefully,
we will change chassis in the next few days.

103: Hotwire, Site tower, Fri 15-Oct-1999 14:36:39 CDT, moving 40m boom
At 1337 CDT, I raised the 40m boom to mount a hotwire, #70. I lowered the
boom at 1342.

102: Status, Site , Fri 15-Oct-1999 11:37:15 CDT, Daily status, Oct 15
ISFF/CASES99 Daily status 15 Oct 1999

Weather: Warm and clear, 10 m/s SSW winds
Surface: dry
IOPs: None scheduled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      X       up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       down 0142 - 0925 CDT, Oct 15
daisy:          up
cosmos:         up
PAM ingest:	intermittent 0142-0500 CDT, down 0500-0945 CDT


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - The 20m sonic died 1040 CDT, Oct 14.  Replaced it with the sonic
   from PAM station 6.  We also moved the KH2O from PAM#6 to 20m.
   The ANL H2O/CO2 sensor at 20m has been removed for offsite repair.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

<><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><>




101: Sonics, Site tower, Fri 15-Oct-1999 10:37:09 CDT, Cycled power on 10m sonic on 60m tower
The 10m sonic on the 60m tower quit around 1755 local time, Oct 14.
Restarted it around 1000 local time, Oct 15.
100: Network, Site , Fri 15-Oct-1999 09:55:50 CDT, marigold jams network
marigold was hung (not pinging, but xmit LED always on) this morning.  We
didn't get data from it starting at 1:42 local this morning.  All the other
adams reported lower data rates but generally stayed up.  The PAM stations
reported data intermittently until 0500.

Marigold rebooted at 0925 and the PAM station ingest restarted at 0945.

99: Kryptons, Site stn6, Thu 14-Oct-1999 18:07:39 CDT, Moved KH2O from Station 6 to 20m on 60m tower
Removed the KH2O from Station 6 at 15:30 local time, Oct 14, and
installed at 20m because we moved the Station 6 sonic to 20m when
the latter died this moring.  Calibrated data not available until 
data system was reconfigured at 17:30.  It was assumed from logbook
entry #19 that this was S/N 1393.  From the calibration sheet in
the case, we used Kw = -0.145, Intercept = 4388, Path = 1.29.


98: Sonics, Site tower, Thu 14-Oct-1999 18:04:22 CDT, 20m sonic died; replaced w/ATI (and KH2O from Station 6)
The ATI at 20m died at 10:40 am local time, Oct 14.  We removed the
ATI and KH2O from Station 6 at 15:30 local time and installed them at 20m.  
The sonic came on line at 16:35.  
97: Hotwire, Site tower, Thu 14-Oct-1999 15:36:16 CDT, Raising 10m and 40m booms to remove hotwires
14:53:00 10m and 40m probes turned to standby in preparation for wx.

14:56:00 CDT raised the 40m boom to dismount the hotwire at 40m. Finished 
at 15:01:00.

15:20:00 raised the 10m boom to dismount that hotwire; finished at 15:24:00.

96: Status, Site , Thu 14-Oct-1999 11:07:46 CDT, Daily status, Oct 14
ISFF/CASES99 Daily status 14 Oct 1999

Weather: Warm and clear, 5 m/s SE winds
Surface: dry
IOPs: IOP6 completed this morning

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD     BAD     up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
 - No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

95: CO2, Site tower, Thu 14-Oct-1999 09:55:16 CDT, No licor cal gases
All of the "lo-cal" CO2 calibration gas was used after the first day that
the Licor was deployed - about 10/7?  I had identified a leak in the solenoid
valve which controlled this line in Tony's box and had tried to fix it, but
apparently was unsuccessful.  There still was plenty of "hi-cal" gas until
sometime during the last 3 days.  Yesterday morning, the "hi-cal" cylinder
also was empty.

Both cylinders are now off.  We probably should disconnect the control circuit
to let the licor sniff real air all the time, but I haven't done this yet.
I don't know how we'll generate a calibration now - perhaps we should manually
attach soda lime to get a zero from time-to-time.  I've put a call in to
Tony.
94: Hotwire, Site tower, Wed 13-Oct-1999 16:45:06 CDT, hotwire replacement/look on 55-m tower (Lundquist)
13 October 1644: Julie raisesboom at 10 m on 55-m tower to replace 10 m hotwire sensor.Julie notes there are a lot of spider webs.Installed probe 69.

1650.  10 m boom is down.

1653:  Julie reports tower vibration; feels like 0.5 Hz.

1701:  Julie is pulling up 40 m boom.Lots of spiders.

1705:  Julie installed probe 70 at 40 m.

170930:  40 m boom is down.










93: Status, Site base, Wed 13-Oct-1999 11:56:06 CDT, Daily Status
ISFF/CASES99 Daily status 13 Oct 1999

Weather: Warm and clear, strong (12m/s) NE winds
Surface: dry
IOPs: IOP6 scheduled for tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD     BAD     up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 00:51 - 11:16 today due to a break (rodents?) in
our fiber-optic cable.  We've replaced this cable now so they have more to
chew on.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- We have rebalanced the AC power load in our trailer and shouldn't suffer
  from further circuit breaker outages.
- We still saw bad signals on stn2 Rpile yesterday, but a cable swap appears
  to have solved it for now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  I tightened a loose
        connection in the CR10 yesterday, but this didn't appear to help.
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan yesterday.	We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 
92: LOG, Site tower, Wed 13-Oct-1999 11:50:43 CDT, tour up 55m tower
Several people were up the 55m tower this morning:
Jielun and Shawn were working on their thercouples,
Roddy and Steve climbed up to 25m for a tour,
Julie worked on hotwires,
Gordon & Tom were replacing fiber cable.
This all between 9-11AM CDT.

91: BaseSystems, Site base, Tue 12-Oct-1999 22:03:23 CDT, writing CDs
We write 4 CDs a day:
	stns			~360 Mbytes
	tower			~350 Mbytes
	film 00:00-12:00	~550 Mbytes
	film 12:00-24:00	~550 Mbytes

	total			~1720 Mbytes/day

The archive hard disk is /data/archive/projects/CASES99/raw_data,
an internal disk on cocklebur, which is 8 Gb.

To write a CD, find a cocklebur window, and
put a new CD into the CD-R drive on top of cocklebur

Execute one of the scripts, passing the month number (10=october, duh)
and the day of the month:

	bkcd.stns mon day
	bkcd.tower mon day

first half day of hotfilm data:
	bkcd.film mon day 0 12
last half day:
	bkcd.film mon day 12 24

This script moves files for that day to the "cdbackup" directory,
and then writes them to the CD.  

IF IT SUCCEEDS, IT DELETES FILES FROM THE HARD DISK, so don't
damage the CD. That's your only archive.

The label is then sent to the printer.
Wait for the manual feed light to appear before inserting a label 
sheet. Slide the label sheet into the manual feed slot on
the back of the printer, until it stops, label side toward you,
with an unprinted label in the lower right hand corner.	

After it prints, peel the label of and put it, sticky side up, 
in the CD Stomper. Put the CD, LABEL SIDE DOWN, DATA SIDE UP!!!! 
over the spindle. Anybody who puts the label on the wrong side of
a CD must stay an extra week!!!!

Put the CD in a sleeve, and stick the narrow label on a strip of
perforated paper and slide it into the top slot of the sleeve (use 
the black plastic thingy).

The log of the CD writing is written to $ASTER/projects/CASES99/cdlogs.

Don't write directly on the CD with any pen - reportedly the ink
can leach through and damage the substrate.








	

	







90: BaseSystems, Site base, Tue 12-Oct-1999 21:47:11 CDT, Hotwire archive
A day's hotwire data is 5.0 Gig.

The hotwire files are on
	/data/blaze/aster/projects/CASES99/raw_data/hw

This is an 18 gig disk, so we can hold 3.5 days.  To avoid
running out of space you should archive yesterday's files
each day.

To archive a day's data to disk:

Log onto clover
Insert a new tape into clover's exabyte, then enter:

	blaze2Tape 0 mon day

where mon and day are the month and day to be written to tape.

blaze2Tape takes about 2 hours or so to complete.  It will also delete the
files of (day-1) if they were previously written to tape.

The blaze2Tape log files and tape listings are on:
	$ASTER/projects/CASES99/blazelists

If you experience tape errors,
	insert a "premium" cleaning tape into the drive
	restart the day's archive by deleting the 2 log files for
	that day on $ASTER/projects/CASES99/blazelists,
	and restarting from the beginning of a tape.




89: TRH, Site tower, Tue 12-Oct-1999 18:04:34 CDT, Replaced hygrothermometer fan at 15m
Sometime this morning, perhaps 10-11 am local time, Oct 12, Steve
replaced the fan in the hygrothermometer shield at 15m.

Oct 14:

Although the temperature error was pretty small, it is now obvious
that there was a definite improvement in the 15m T following this
replacement.

Unfortunately, RH.15m is still about 2 %RH lower than RH.5m and RH.25m
during the day.  At night RH.15m is lower than both RH.25m and RH.35m.
The 15m mixing ratio is also anomolously low at all times.
88: Radiation, Site stn2, Tue 12-Oct-1999 17:42:32 CDT, Rpile.out at Station 2
The Rpile.out data has been intermittent at station 2.  At 3:50 pm,
local time, Oct 12, I disconnected and reconnected the cable at the
sensor and this appeared to solve the problem.  At that time, I also
bumped the Schwartz out of level by popping the leveling screws out
of their upper seats.  I reseated the screws and found that the level
was still okay.  At that time I also adjusted the level of the net
radiometer about its N/S axis.

The Rpile.out data soon failed again and at 5:34 I swapped the pile.in
and pile.out cables at both ends.  This fixed the problem again.

87: ADAMS, Site base, Tue 12-Oct-1999 16:49:47 CDT, How to restart adams & eves
Bring up the console window on ASTER so you can watch the adam/eve bootup
messages.

	ping cosmos

If it is alive, you can probably bring it up with mxreset:

	mxreset cosmos

Otherwise you will have to power cycle it.	

In the console window you should see: 

Oct 12 16:52:03 cosmos  System reset!
Oct 12 16:52:29 cosmos  adatsSock SO_SNDBUF len=4096
Oct 12 16:52:29 cosmos  adatsSock SO_SNlast message repeated 1 time
Oct 12 16:52:30 aster cosmos[753]: cosmos Beginning connection on port 1030
Oct 12 16:52:31 cosmos  cosmos booted
Oct 12 16:52:35 aster cosmos[753]: cosmos STREAM socket connected to aster:33008, 1 active connections
etc...etc...

Then do the same for daisy,marigold,ragwort.

If an adam is pingable, but mxreset does not work from ASTER, then try:

	rlogin cosmos
	VxWorks login: aster
	Password: asterize
	-> mxreset

To restart the PAM ingest processes, log into toucan, and do:
	sio_shutdown -s

This should kill 'em all.  If curious, check:
	ps -ef | fgrep stn

If things are OK, then restart them with:
	serial_ingest -z




86: Sonics, Site stn4, Tue 12-Oct-1999 15:08:17 CDT, Tilt of PAM sonics and mean w
The mean vertical velocity for PAM station #4 appears anomolous 
compared to the other PAM stations.  The explanation is apparent
from tilt plots.  The sonic at station #4 is tilted about 180
degrees different from those at the other PAM stations.
85: Hotfilm, Site tower, Tue 12-Oct-1999 14:06:10 CDT,
9 Oct  1100 to 1700 Tried to find spiking problem.
9 Oct 1715  Steve/Gordon changed timing of zero circuit relative to sampling of sensors,  seems to take care of spiking problem. 
After this time the zero circuit is left on.
Before this time the zero circuit was off and base voltage was left to drift.
It was reset most days by re-installing circuit for short time. 
Spectra show spikes at 60 and 80 hz in the two cold films but not in the hot films. Spikes are small - we will probably remove them from the data after the fact.
10 Oct - 1145 to 1240 - reset timer on zero sheath covers on hotfilms.  Timer has been set wrong at zero 0300 and 1100 and 1200 hours.  Now set to cover sensors at hh:59 each hour.


84: LOG, Site , Tue 12-Oct-1999 11:11:28 CDT, notes for Steve Semmer
In preparation for our shift change in a few days, this is a list of changes
from the set-up configuration:
- I've labeled 3 T/RH fans as dead, broken, and weak.  We still have more spares
- Various people are using our tools.  I've pretty much let them, and most
  have come back.
- The Dutch have been charging their battery in the back of our trailer.
- Of course, the Dutch and Spanish people have towers, stakes, etc., which
should come back to us.  The boom clamps used by the Spanish are ours. 

Spare status (10/13):
- Our only sonic is acting weird.  I intend to bring it back to ATI with me.
- We haven't tapped the T/RH spares yet, but I'm thinking about it for the
  15m sensor.
- One prop is bad and is now (labeled) in the bottom of the prop packing box.
  There is another spare in this box.
- We've just used our "spare" 300m fiber cable to replace the other one
  which is now chewed in 3 places (but still has one good fiber).
- We have no spare EVE components.
- We had 2 spare Tsfc probes, but one is labeled "bad" now.  The other is
  being used by Jielun for manual Tsfc mapping.  Grab it from her if needed.
- fester/lurch has been intermittent and probably is not available as a
  spare adam.  The chassis should work (and should be used to replace
  ragwort's sometime).  There are lots of spare ADAM boards, but I don't
  think any work in combo with others -- check with Gordon.
- The Toshiba is the best laptop!  The Compaq has been retired since it has
  a bad serial port.  Rocky is still available, but its battery only lasts
  about 10min and the 12 VDC adaptor doesn't work.  The Ambro is available,
  but only Gordon knows how to get it to boot!
83: Status, Site base, Tue 12-Oct-1999 10:04:38 CDT, Daily Status
ISFF/CASES99 Daily status 12 Oct 1999

Weather: Warm and clear, S winds at ~5 m/s
Surface: dew this morning, dry by now
IOPs: IOP5 finished this morning (5pm - 8am); none tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD?    BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up; failed twice yesterday (11:35-13:05 & 14:10-18:25 CDT) 
		due to overheating
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 07:50-08:50 due to a blown circuit breaker.
The blown breaker also caused lost data from the 10m towers 08:15-08:30.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- An overloaded circuit in our trailer caused power outages through the night.
  We will map out the circuits today to fix this.
- Resolution changed on all 10m kryptons, so we are now able to see smaller
  fluctutations.  We changed to "ops2" in the beginning of IOP5.
- The sonic at 55m was dead through this IOP.  We cycled power this morning
  which revived it.  (We should have done this last night.)
- I don't see noise on the stn6 krypton now.  This may be due to changing
  this station's config not to sample pp.
- We saw a bad signal on stn2 Rpile yesterday, but it appears to be ok now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  Dave Auble found some questionable connections which he
	tightened yesterday.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating and are now connected to
	EVEs at stns 1 and 5.  He hasn't provided a cable to connect to
	stn4 yet.  Our resolution on this signal is about 5 ubars, which may
	be too coarse in some cases (but this is the best we can do).

82: Sonics, Site tower, Tue 12-Oct-1999 10:02:48 CDT, ati.55m power cycle
I cycled power on 55m ati which brought it back to life.  This done about
9:30 (and I was up the tower to ragwort).  This sonic has been down since
2pm yesterday - oops...
81: ADAMS, Site tower, Mon 11-Oct-1999 21:25:41 CDT, new Ironics card at ragwort

Installed new Ironics 3234 card at ragwort. pp.50m shows spikes due
to dropped characters, which I assume is due to the ironics 1624
not keeping up.



80: ADAMS, Site tower, Mon 11-Oct-1999 19:06:55 CDT, ragwort overheating
Ragwort died today at 11:34 CDT (16:34 GMT).

The LMON light on ragwort's port in the chem shelter fiber box 
was out.  Before I realized it was probably a cooling problem,
I just switched the breaker on and off.

Ragwort came up but
	siostatus yarrow
showed non-zero values in the msg ERRORS column.

I then did a mxreset and it came up with the serial card working.
It then died at 19:08 GMT.

At 23:00 I climbed the tower.  The power supply was dead - the
three red lights were out.

I redid the ventilation cardboard so that the air flows over the
power supply in addition to the 4 VME cards (but not the signal-cond
card).  Also ripped out the insulation under the top of the adam.

Finally looked at Tbox.ragwort.  A plot of Tbox.ragwort shows that
the cooler must have failed at about 06:40 CDT this morning.
Tbox went from its 10 minute cycling at about 27 degC up to 35 degC 
at 11:30 when ragwort failed.

After restarting ragwort at 18:30 Tbox shot up to 34C, so at 20:00
the lid was left propped open about 5 inches.







79: Kryptons, Site , Mon 11-Oct-1999 18:51:22 CDT, all stations have new krypton gain
I've just loaded new krypton.dat files into all stns.  This changes the
krypton voltage range to +/-1V, which hopefully will be enough, but also
increases our resolution by a factor of 10 to ~0.5mv.  

This change also (finally) configured stn1 to have pp and not stn6.

We will have to change ops (during an IOP!) to accomodate this change
in the kryptons.  I'll do that now.
78: CO2, Site tower, Mon 11-Oct-1999 16:33:54 CDT, 20m H2O/CO2 reinstalled
We just put the ATDD sensor back up.  Dave Auble simply tightened connectors
inside the electronics/power supply box and removed two passive filters
added by ANL.  The data look almost believable now.

77: CO2, Site tower, Mon 11-Oct-1999 13:26:40 CDT, 20m boom up/CO2 removed
We've just removed the 20m H2O/CO2 sensor so that Dave Auble can look at
it today.  This was from about 12:30-13:00
76: Status, Site base, Mon 11-Oct-1999 11:12:47 CDT, Daily Status
ISFF/CASES99 Daily status 11 Oct 1999

Weather: Warm and clear, SW winds at 5 m/s
Surface: dry
IOPs: ran early this morning (4am - 8am) and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      noisy   -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We cleaned all kryptons to bring up the signal levels 
  and may change the A/D gain.
- We replaced the sonic at 10m

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  Cleaning did not change this.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.





75: Kryptons, Site tower, Mon 11-Oct-1999 11:08:28 CDT, Cleaned 5m KH2O
Cleaned 5m krypton hygrometer windows between 10:45 and 10:50 local time,
October 11.


74: LOG, Site , Sun 10-Oct-1999 19:41:21 CDT, Nappo pp connected to stns 1&5
I've added connectors to the ends of cables provided by Dave Auble.
However, stn1 still doesn't have a config for this signal - I'll do
this tomorrow.  (I had problems with rocky, including a blown fuse, tonight.)

Auble hasn't provided me with a cable out of stn4 yet.

73: Sonics, Site tower, Sun 10-Oct-1999 17:09:41 CDT, 20m ATI was at 10 samp/s
Somehow, the 20m ATI was set to 10 samp/s.  I assume that this has been the
case since set-up.  It is now at 20 samp/s.

72: Sonics, Site tower, Sun 10-Oct-1999 17:05:40 CDT, replaced 10m ATI
Something has been wrong with the 10m ATI since it was installed.  The worst
signal is just w.  I just replaced it with our "spare" which was removed from
40m earlier due to spiking and never changed.  I'm hoping that this will
work ok in the lower wind speeds at 10m.  (It ran on the bench just fine.)

Note that this entailed switching the hotwire mount, which fortunately was
straightforward.

Also note that this sonic also has a level, so the config is the same.
(Unlike when we swapped at 40m -- the sonic now there is reporting dummy
values for the levels (i.e. analog output is turned on, but there is no
signal.))

71: Kryptons, Site stn4, Sun 10-Oct-1999 14:28:45 CDT, cleaned stn 1-6 kryptons
I cleaned all kryptons today, which brought up the signals significantly.
All were cleaned by lowering the masts.  I disconnected the krypton signal
while the mast was down, which should indicate when I did this - approximately
10-1pm today.  I worked on the stations in the order: 6, 4, 1, 3, 5, 2.

The sonics at 1 and 5 were loose on the booms, so I tightened the screws.
70: Sonics, Site tower, Sun 10-Oct-1999 10:49:33 CDT, 20m boom orientation
I rechecked the 20m boom this morning and found that the mount had
shifted to the south, presumably caused by the process of pushing the
boom to the north when raising it (to avoid crossed guy wires above it).

At 10:20 am local time, Oct 10, I moved the mount to its correct 
position that was marked by black electrical tape.  Unfortunately, 
the boom still appears to be angled a few degrees to the north 
relative to the other booms.
69: Status, Site base, Sun 10-Oct-1999 08:45:09 CDT, Daily Status
ISFF/CASES99 Daily status 10 Oct 1999

Weather: Warm, some low-level fog, clear above, light winds
Surface: dew
IOPs: ran last night and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) WET     -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      ok      -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We'll clean these to bring up the signal levels and may
  change the A/D gain.
- We replaced the T/RH fan at stn3.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  We'll clean it today.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  In today's case,
	fog has made the 1.5m sensor bad, but others are okay.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
68: Status, Site , Sat 09-Oct-1999 23:05:17 CDT, evening IOP status
An IOP is now in progress, but we are going home (it is 11pm and we've been
here since before 9am).  I've just scanned all cockpits, and the only obviously
bad sensors are the 20m H2O/CO2 sensor and the stn1 Gsoil.  Hopefully, all
will stay going.

I also noticed a thick haze layer at sunset which should be good for radiative
flux divergence.
67: LOG, Site tower, Sat 09-Oct-1999 22:32:57 CDT, 5m tower down times
5m tower down:
Oct 1 most of the day
Oct 3 tower down 9 to 1100.
Releveled tower and sensors at 1000
Oct 5 1630 - 1804 tower down. Replaced 5m cold film
Oct 5 1900 - 1920  grounded cf bridges to bridge power supply



66: Kryptons, Site stn1, Sat 09-Oct-1999 21:36:42 CDT, looked at krypton voltage
I just disconnected the krypton from stn1 for a few minutes while checking
its voltage level (quite low: 0.090 V).  It is reconnected now.  This 
done about 2115 local.
  
65: Kryptons, Site stn3, Sat 09-Oct-1999 16:01:51 CDT, cleaned krypton on stn3
I just reached up to clean the krypton on stn3 about 10 min. ago.  It may
take a few minutes for this sensor to recover.

This done because Tom had noticed low fluxes on this sensor.
64: TRH, Site stn3, Sat 09-Oct-1999 15:59:51 CDT, T/RH needed new fan
The fan had died on stn3 T/RH.  An hour ago, I pulled the sensor off this
station, replaced the fan with a new one, and just reinstalled the sensor.

Tom thinks this has been bad since at least 29 Sep.


63: LOG, Site , Sat 09-Oct-1999 12:41:31 CDT, maintainance tasks
Now that we are operational, here are some of the regular tasks that should
be done:

Daily:
- status message to gstoss@ucar.edu, cole, dcarlson (and others in Surf Grp)
- archive data (more on this in a later entry)

Weekly:
- clean kryptons with distilled water and Q-tips
- clean pyrgeometer domes? (including "boom"?), with breath and Kim-wipes
  (I am a bit worried about the hauling line breaking)
- soil moisture checks at stns 1, 2, 3, & 5
- tower cable tension measurement - examine anchors
62: Tower, Site tower, Sat 09-Oct-1999 12:17:31 CDT, raised and lowered 20m boom
From 10:45 to 11:25, the boom at 20 m was raised and lowered so that we could 
switch out hot wire probes (took out #78, put in #71) at that level. 
Inspection showed that #78 broke because a tower denizen spun its web across 
the probe, probably incurring quite a jolt. 

When the boom was lowered, it appeared that the 20m boom was angling off to 
the north more than the other booms. Tightening the south guy wire did not 
change that.
61: Status, Site base, Sat 09-Oct-1999 11:21:29 CDT, Daily Status
ISFF/CASES99 Daily status 9 Oct 1999

Weather: Warm, 2/8 fair-weather Cu, light winds
Surface: dry
IOPs: scheduled for tonight and tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      BAD
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 

General:
- We're finally operational!  Some sensors are not operating properly, but
  most critical sensors should provide good data.  We are ready for this
  IOP.
- daisy now appears to be working
- We replaced the 55m krypton hygrometer and 15m prop which fixed problems
  with these measurements.
- Spikes on CSAT sonic anemometers did disappear when they dried off.
- Spikes on some ATI sonic anemometers may reappear, but probably not in
  light winds.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.
     - The fan on the T/RH sensor at stn3 may not be working.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
60: PropVanes, Site tower, Sat 09-Oct-1999 10:58:27 CDT, swapped prop at 15m
The prop at 15m (W0002) has been reading nearly 0 speeds for quite a while.
I just spent the last hour+ swapping it with spare sensor (W0003).  I
initialized this spare in the trailer with new propeller P56159 with
a pitch of 306, reset the boom alignment (only changed by one count from
what it had initially), and set the boom angle to 97.0 deg.  (I note that
W0002 had 96.7 deg, which is off by 0.3 deg from the boom angles which
were logged.)  I put the old vane in the new sensor.  It bent slightly when
I pulled it into the tower (even when I was watching for this), but I think
I bent it back close to straight.

The speed profile appears to be better now.

59: CO2, Site tower, Fri 08-Oct-1999 17:07:37 CDT, co2.20m fixed itself?
co2.20m appears to have a reasonable signal now, despite no action being
taken on our part.  It definitely has been offscale or spiking for the
previous several days.  We'll keep watching it.

58: Sonics, Site tower, Fri 08-Oct-1999 16:50:42 CDT, Moving the 40 m boom
This entry should have been made on Tuesday, October 5. From 12:40-13:03 CDT, I raised and lowered the 40 m boom to mount a hot-wire probe next to the sonic.

57: Kryptons, Site tower, Fri 08-Oct-1999 16:21:01 CDT, new krypton at 55m
By swapping things, we found that the head of S/N 1133 is responsible for
the 50s noise (and generally bad signal) from h2o.55m.  We've swapped the
whole sensor with S/N 1389.  We'll update cal_files soon.

P.S. Obviously, we (Tom and I) were up the tower -- from about 3:00-4:10.

56: Radiation, Site stn1, Fri 08-Oct-1999 12:32:09 CDT, Corrections for radiometer deployment errors
I inventoried the surface radiometers to investigate the cause for
some anomolous long wave data:

Station	Rnet	Pyg.in	Pyg.out	Psp.in	Psp.out
 1	97034	29260	26416
 2	94197	31977	31981	940181	940187
 3	94125		31975
 5	97035		31979

This showed that the radiometer arrays on stations 1 and 2 were
interchanged when deploying them at the end of the intercomparison
period.  This morning (Oct 8) we entered into the CR10 loggers at 
stations 1 and 2 the coefficients corresponding to the radiometers 
mounted there.

WARNING  WARNING  WARNING  WARNING  WARNING  WARNING 
Be aware that the labeling of the radiometers at stations 1 and 2 
is different for the intercomparison than it is after the radiometers 
were deployed at the various PAM stations.
WARNING  WARNING  WARNING  WARNING  WARNING  WARNING 

The coefficients for the PSPs at station 2 were entered around 9:30
am local time.  
The coefficients for the PYGs at station 1 were entered around 10:55
am local time.
The coefficients for the PYGs at station 2 were entered around 11:10
am local time.  At that time we found that the signal cables for the
two PSPs were interchanged and this was corrected at the CR10.

At station 5 we found a small error in the third coefficient of the
Tcase.pyg.out calibration.  The data sheet lists a coefficient of
-6.90295 and the CR10 has a coefficient of -6.9229.  It was decided
to leave this coefficient as we found it, since this value was used
during the field intercomparison of the radiometers.

55: Status, Site base, Fri 08-Oct-1999 11:47:08 CDT, Daily Status
ISFF/CASES99 Daily status 8 Oct 1999

Weather: Warm, Overcast, light winds, hint of drizzle
Surface: damp
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      BAD     -       -       -       -       -
50m(r)  WET     -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  SPIKES  ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  WET     -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          DOWN - NO DATA FROM 15-25m
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Data appear to be unreasonable.  We will check and/or replace hygrometers
	at 55m and stn6 today.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds -- we may fix our spare and
	replace this sensor.
WET   - Campbell sonics often fail with moisture on the transducers.  The higher
	sensors have been in fog all morning, but should resume good operation
	once dry.  

DOWN  - Daisy has recurring problem with crashes -- we are working on this.



54: CO2, Site tower, Fri 08-Oct-1999 09:03:30 CDT, approximate licor calibration
A crude fit between the Licor and co2.40m gives:

CO2 = (Licor + 2.4)*140

in mg/m3 of CO2.  To convert to ppm:

mg/m3 * g/1000mg * mole/44g * 22.4 lit/mole * m3/10^3 lit * 10^6ppm ~ 0.5

With these factors, the cal-gas values are about:
357ppm; 330ppm; 364ppm

I will enter these values in prep.config.


53: Radiation, Site , Fri 08-Oct-1999 08:52:04 CDT, Rnet cal factors
Rnet calibrations from REBS 8/17/99 calibration are:
	Positive Negative (Zero Wind CF in W/m2 mV)
94196	9.25	11.44
97035	9.28	11.54	stn5 (stn5 is using these values)
94197	9.07	11.25	(stn2 is using these values)
91096	8.54	10.59
94125	8.65	10.75	(stn3 is using these values)
97034	9.31	11.39	stn1 (stn1 is using these values)



52: CO2, Site tower, Fri 08-Oct-1999 08:44:28 CDT, Licor co2 installed
Yesterday, I found that the Licor problem was wrong pins connected to the
CO2 analog output channel and corrected this problem about noon.  Signal + 
was connected to power ground (pin7) and Signal return to open (pin4).  
The signal now appears to be working.  At that time, I also opened the cal
gas cylinders and installed the valve control lines.  I see 2 calibrations
in the data at about 4PM and 4AM last night.


51: ADAMS, Site tower, Thu 07-Oct-1999 17:13:10 CDT, daisy problems
Daisy is sick.

Configuration:

  daisy
  chicory
  datel
  chassis #4

Did lots of switching cards around.  

Indications also pointed to cooling problems, because the adam would 
work fine and then die 1 half hour after closing the lid.
(Daisy's lid is missing insulation, which should keep it cooler.)

The new matrix card has its own problems.  After a while it would
choke on a bad sample (chan==0,length==0).  This appears to be
some sort of memory corruption problem.

The new ironics cards may work now, but perhaps run hot.
They have 4 meg of RAM (the others have 1 meg).

Ventilation was improved by duct taping cardboard in the fan "horn"
with gaps only at the ends so that the cool air is directed only
to the 4 cards on the left (matrix,ironics,serial breakout and datel)
and to the power supply on the right.

Here is what is working:

  daisy matrix (S/N #9010
  chicory ironics (#014029)
  daisy's original datel (#440690)
  PSU power supply
  chassis #4


50: Sonics, Site , Thu 07-Oct-1999 16:32:50 CDT, ATI sonic gain
Can change R42 on first stage amplifier.  Default value is 33.2K, could
change to 49.9K to obtain higher gain.
49: Radiation, Site , Wed 06-Oct-1999 17:07:04 CDT, radiometer work
Stuff done today:
stn1: changed CR10 baud rate to 1200.  Commands are:
	*0 
	7H 
	*1 
	A (until step 62)
	51 (overwrites 52)
	*0
	(this can be done through EVE by typing:
	entersys	
	stop
	talk (port 4, rate 9600, etc.)
	start
	Rnet is Q97034
stn2: moved solar panels; hooked up battery swiped from stn5; mounted Rnet
	with 2 screws (only had one before); leveled Rnet; added connector
	to bare wires from Rnet so it could plug in; changed baud rate to 1200.
	Rnet is labeled "station2"
stn3: moved solar panels; moved rad stand from IC array; leveled and oriented
	stand; mounted Rnet from box labeled "station3", but with a label of
	"station1"!  Found that fan cable wires were color-coded backwards, so
	rebuilt this cable (red negative, black positive); changed baud rate
	to 1200.
stn5: stole battery from main station to power radiometers at stn2; removed 
	wrong DC power cable; attached good DC power cable; cut cable ties;
	changed baud rate to 1200.
	Rnet is Q97035.
All rad stand solar panels are now located about 4m to the NE of the CR10 box,
	with the CR10 on the East side of the rad stand.
48: CO2, Site tower, Tue 05-Oct-1999 19:41:05 CDT, co2 turned off
turned off Licor (includeing pump and cal gases), since it has been reporting
bad values.  This done about 6pm local.  It has been running bad for the last
day and a half.

47: Radiation, Site , Tue 05-Oct-1999 19:38:10 CDT, moved rads out of IC
Installed 48m boom about 1300 today - leveled to within 0.003 deg.  The
pull rope is tight!  Also, don't trip over the CR10 box on the floor 
at 49 m.

Mounted radiometers at stns 5, 2, and 1 (in that order), but not reporting yet.
stn5 is down waiting for a power cable/solar power move and baud rate change.
stn1 is waiting for a baud rate change
stn2 is waiting for a new battery!
stn3 is waiting for the whole rad stand to be moved.
46: Network, Site stn4, Mon 04-Oct-1999 14:42:38 CDT, communication up to stn4
stn4 showed signs of bad communcation, since some bytes were lost.  This has
been going on for many days.

I replaced the max487 chip on the 485 converter board, which appears to
have brought the station back.  (Thanks for sending these, Steve S.!)

45: TRH, Site stn2, Mon 04-Oct-1999 14:40:36 CDT, fix of T/RH problem?
I switched cables (5m instead of 10m to the T/RH) and the Y-adapter on this
station and appear to have gotten both the T/RH and prop to work.  It appears
that the T/RH cannot drive 20m of cable with lots of connections.

In the process, I swapped T/RHs between stn5 and stn2, so ID=10 is now at
stn5 and ID=12 is at stn2.

44: Sonics, Site tower, Mon 04-Oct-1999 10:53:32 CDT, Movement of booms
From 0930 to 1045, Horst and Lundquist raised and lowered the beams at 
40m (to remove a hotwire probe shield and support), 20 m (to mount a probe),
and 10 m (to switch out shield/support and mount a probe), in that order.

43: Status, Site base, Sun 03-Oct-1999 11:50:32 CDT, Daily Status
ISFF/CASES99 Daily status 3 Oct 1999

Weather: Cool, Overcast, lighter winds
Surface: reasonably dry
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  up      up      up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      up      up      up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
1.5m(c) up      -       -       up      -       -       pp up

Data Systems:
ragwort:        DOWN - NO DATA ABOVE 40m YET
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC/Nap] up     up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       up      up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  H2O/CO2 sensors at 20m 
	and 40m now mounted and should be connected soon. Licor CO2 at 1.5m
	may be installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       Ragwort has recurring problem with crashes.
42: Status, Site base, Sat 02-Oct-1999 09:34:16 CDT, Daily Status
ISFF/CASES99 Daily status 2 Oct 1999

Weather: Cool, Overcast, windy
Surface: reasonably dry - might have been some rain overnight (but not
	before we left at 23:30)
IOPs: none, since rain predicted through the weekend.

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      ready   ready   up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
2.5m(c) DOWN    -       -       up      -       -       pp up

ragwort:        DOWN - we suspect that heating is the problem
marigold:       up
daisy:          up - back to normal (though still some crashes)
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  Will try to get the CO2
	sensors installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at 4 and 5 only.  Their data system will be at 1
	(but we won't get data from it).
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       csat.1.5m didn't come up after the software upgrade.  We need to figure
	out how to configure it properly.
       Ragwort has recurring problem with crashes.  We're trying to improve
	cooling (either by removing insulation or resetting the thermostat)
	to the box -- also for daisy.

41: Radiation, Site , Fri 01-Oct-1999 23:04:50 CDT, cleaned domes
Just cleaned domes on all pygs in the IC array.  They all were somewhat
dirty.  Also checked levels and all were good.
40: Status, Site base, Fri 01-Oct-1999 08:33:21 CDT, Daily Status
ISFF/CASES99 Daily status 1 Oct 1999

Weather: nice.  Clear, not too windy
Surface: reasonably dry
IOPs: none, since rain predicted to start tomorrow through the weekend,
and tonight should be windy

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   [Mill]  up      ready   up      -       up      -
2.5m(c) [Mill]  -       -       up      -       -       pp up

ragwort:        DOWN - has daisy's parts.  Being worked on today.
marigold:       up
daisy:          DOWN
cosmos:         up. Was down last night by accident


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      up      DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	I've started looking at data, but need to do this more.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send one H2O/CO2 sensor.
       The other is in hand, but is missing an AC power cable (but we may just
       plug this into DC anyway)
[Mill] - Dave Miller (Univ. Conn.) is set up except for his 1.5m sonic.  This
       was awaiting a part which arrived yesterday, so should be up today.
[Nap] - Carmen Nappo (NOAA/ATDD) did most installation yesterday - we need
       to be sure that these sensors are cabled into EVE.  Should be up today.

DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.
       Cosmos was left switched off last night.  It is up now.
       Ragwort has recurring problem with crashes.  We're in the middle of
       board swaps to fix.

39: ADAMS, Site tower, Thu 30-Sep-1999 12:30:00 CDT, ADAM status
Swapped the matrix board in ragwort with daisy's.

Results:

  daisy could not talk to datel (HSTSTA errors)

  ragwort is sick, will not stay up. 
  After a while it does not see ironics board and mxreset does not
  cure it. A power cycle will work for a while

Then on 10/1
  put the daisy matrix board back in daisy:  works

  put the PSU matrix board into ragwort.  The EPROM is configured
  for 1 meg RAM without PPP.
  
  The PSU ragwort booted, but then "siostatus yarrow" showed msg errors.
  After a while it died.  Same symptoms as the ragwort matrix.


   

38: ADAMS, Site tower, Wed 29-Sep-1999 21:07:06 CDT, ADAM status

Power to tower was down last night due to a rabbit's chewing.   

I was distracted so didn't monitor adams much today. After power
was restored the adams were up except for ragwort.

Ragwort was dead (status light orange, Ironics halted). 

Replaced the Ironics 1624 in ragwort (yarrow) with a new iv3234.
37: Kryptons, Site stn3, Wed 29-Sep-1999 16:01:29 CDT, Krypton spiking problem w/EVE A/D
Krypton data from sensor 1391 was glitching as shown on the
cockpit display.  It was the only sensor showing this trend.
A check running 'adflag' of the station revealed apparent
dropped a/d counts on the first cycle of the 20hz krypton task
after a reading taken of the .5hz power monitor parameters.
Approx 50 counts would be missed.
This is symptomatic of the a/d problem usually experienced
by the power monitor which can have a reduced vbatt reading.

Station 3 electronics box has been the poorest performing
PAM station in terms of responsiveness with the fastout and
high speed ingest processing.  A couple of vx errors were
observed.  This is probably a timing problem between the two
analog tasks and the device driver.   It is likely either
the a/d board, dcom332 cpu or the backplane or a combination.

A/D board swap to station 1
From station 3 board 439220 removed and swapped into sta 1
From station 1 board 439176 removed and put in sta 3.
This should reveal whether the glitching problem follows the
A/D board.  I don't think it will.
36: Radiation, Site stn3, Wed 29-Sep-1999 15:58:52 CDT, Tsfc replaced on PAM-3
Everest IRT Swapped:
Sensor 3577-4 removed; sensor 3577-2 installed

After the thunder/rain storm 2 nights ago the old sensor
was reading 100 deg-C.
35: PAM-EVE, Site base, Wed 29-Sep-1999 13:10:10 CDT, PAM station Configuration Summary
The following configurations are available on ASTER
in /home/pam/config and individual station-ready versions
in /home/pam/projects/CASES99/sta_cfg.

On the Rocky they are located in:
c:\pam\cases99\fmt0070
c:\pam\cases99\fmt0071
c:\pam\cases99\fmt0072
c:\pam\cases99\fmt0073
c:\pam\cases99\fmt0074
c:\pam\cases99\fmt0075
c:\pam\cases99\fmt0076

Feature		70	71	72	73	74	75	76
FASTOUT		y	y	y	y	y	y	y
Sync10,despike	y
File metppf	y	y	y	y	y
File /card	y	y	y
File 'cardcopy'				y	y
GPS, timesync	y	y	y	y	y	y	y
stations		1-3,5	4,6	1-3,5	4,6	1-3,5	4,6
logger		y	y		y		y
ATI		y	y	y	y	y	y	y
SBUS		y	y	y	y	y	y	y
power		y	y	y	y	y	y	y
ptb220		y	y	y	y	y	y	y
tsfc		y	y	y	y	y	y	y
< krypton.dat	y	y	y	y	y	y	y


Summary:
	fmt0070	- Used during testing only, not for ops.
		includes sync10 and despike and handy for attempts
		to work on reboot eve code because with 20hz ati,
		a/d and fastout this is good for locking up some 
		stations.

	fmt0071 - Initial implementation includes data storage to
		both ramdisk and /card
	fmt0072 - Same as 71, except the logger is not included.
		use for stations 4,6

	fmt0073 - Primary Config, stations 1,2,3,5
		Basically same as 71 except uses 'cardcopy' feature
		available with eve prom 990926 which does a 1/day 
		card copy instead of writing to it every 5-minutes
	fmt0074 - Primary Config, stations 4,6
		Same as 73 except no logger.

	fmt0075 - Backup contingency if EVE slogs down.
		Minimal config. does not have local data storage.
		IE Do Not Use this if the fastout port to aster and
		archiving is questionable, down, etc.!!
	fmt0076 - same as 75 but no logger

Krypton Declaration Files:
	Two different PAM config files for each krypton hygrometer have
	been created.  One each for sampling with the delta-pressure sensor
	and one for sampling the krypton alone.  The calibrations are
	based on data from Campbell full range, normal conditions.
	The files are called 'kryp13xx.dat' for sensors 1389-1395
	Declaration files are found in directories:
	Without Delta Pressure: for stations 1,2,3:
		On Rocky c:\pam\cases99\krypton\no_dp\
		On aster /home/pam/projects/CASES99/sta_cfg/krypton/no_dp
	With Delta P included for stations 4,5,6
		On Rocky c:\pam\cases99\krypton\with_dp\
		On aster /home/pam/projects/CASES99/sta_cfg/krypton/with_dp


34: PAM-EVE, Site , Wed 29-Sep-1999 11:46:01 CDT, PAM FastOut Port Outages due to EVE-RS485 driver
PAM Stations 5 and 6 had their FASTOUT connection to Toucan/Aster
interrupted after a heavy thunder/rain storm 2 nights ago.
In both cases the problem was not the cable but the EVE RS232/485
converter module located on the EVE electronics interface front panel.
The 8-pin socketed MAX487 chips on the module were damaged apparently
by a common-mode ground surge.  Replacement of these chips fixed the problem.
In one case the module was swapped with the identical module used for the
sbus and the damaged unit was still able to drive that short mast cable at
least.


33: Radiation, Site , Tue 28-Sep-1999 18:08:51 CDT, playing with rad biases from IC
Here is a first cut at adjusting biases on Rpile from the radiometers
based on data from the interval "99 Sep 25 17:00",lenhr=31,stns=0:

Rpile.in.a.boom Rpile.in.b.boom Rpile.in.s01 Rpile.in.s02 Rpile.out.a.boom 
              0             0.5         -1.2          bad             0.7
Rpile.out.b.boom Rpile.out.s01 Rpile.out.s02 Rpile.out.s03 Rpile.out.s05 
             0.4            -1           bad           0.2            10

32: ADAMS, Site tower, Tue 28-Sep-1999 09:15:38 CDT, ADAM status
Lots of rain last night with lightning.  Must have had power outages
because the adams all rebooted at the same times.  Toucan reboots
cause breaks in the PAM archive.

The log from the UPS in the trailer showed some entries for these
times also.

01:01:50	cosmos  ?? (probably not a power outage)
01:55:00	cosmos,daisy,marigold,ragwort,toucan
02:20:30	cosmsos,daisy,marigold, toucan
02:21:30	cosmos,daisy,marigold,ragwort
03:10:50	cosmos,daisy,marigold,ragwort,toucan
12:01:30	cosmos,daisy,marigold,ragwort,toucan

Everything was running in the morning, except the serial channels
on ragwort.  "siostatus yarrow" showed errors in the "msgs" column.
mxreset would not bring the serial channels back.  Power cycling
ragwort from the base of the tower did the trick.

The last network fix to ragwort was swapping the fiber transceivers
with cosmos.  That must have helped because we haven't had any
unexplained reboots since then. 











31: Status, Site base, Mon 27-Sep-1999 17:17:41 CDT, Daily Status
ISFF/CASES99 Daily status 27 Sept 1999

Weather: windy, drizzle, cool all day (Severe T-storms last night)
Surface: plants wet, roads muddy, ground still hard
(lots of visitors today: Hatfield, Miller, Ned & Tim, etc.)

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  DOWN    [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   [Mill]  up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        up
marigold:       up
daisy:          up
cosmos:         up-one crash


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Dave Miller (Univ. Conn.) is now on-site and working on his
	sensor mountings -- a chem shelter has been emptied and moved for him.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.

DOWN - 40m ATI was spiking yesterday, so we removed it.  We swapped stn4's
	ATI with our spare UW to have another ATI to swap back in.  I added
	the hotwire mount to this, but both ATIs are now in the lab.  (It was
	too wet to climb today.)


30: Radiation, Site , Mon 27-Sep-1999 11:55:30 CDT, releveled radiometers
I just used the small bubble level to check the levels.  Most looked okay,
though I adjusted the western 2 and the eastern 1 a bit.

29: ADAMS, Site tower, Sun 26-Sep-1999 08:54:50 CDT, ADAM status
cosmos:
  ran overnight
daisy: 
  ran overnight
marigold:
  ran overnight
ragwort:
  dead this morning.
  From archive:
     died at 06:18 GMT (01:18 am CDT)
     restarted itself at 06:25 GMT.
     died at 08:30 GMT
     restarted itself at 11:43 died immediately
     rebooted at 14:43 (9:43) by human

  Log file on toshiba was empty - serial
  connector fell off after setting it up.

  Swapped fiber transceiver with the one from cosmos.  Rebooted at
  15:47 GMT

  








28: Sonics, Site tower, Sat 25-Sep-1999 16:43:03 CDT, sonic parameters
adam	chan	height	type	owner	serial

cosmos	200	2.5m	CSAT	Miller	
cosmos	201	5m	CSAT	Miller
cosmos	204	10m	ATI	NCAR	980505 (w/levels)
cosmos	205	20m	ATI	NCAR	980202
marigold 200	30m	CSAT	Mahrt
ragwort	200	40m	ATI	NCAR	980504 (w/levels)
ragwort	201	50m	CSAT	Mahrt
ragwort	202	55m	ATI	PNL	980415

All ATIs:
Sample Options Menu\r\n
A. Median Filter (Off)\r\n
B. Data Quality Algorithm (Off)\r\n
C. Shadow Correction Algorithm (Off)\r\n
D. Remove RH from Temp Calculation (Off)\r\n

27: ADAMS, Site tower, Sat 25-Sep-1999 16:20:53 CDT, ADAM status, turned off chem shelter AC
cosmos:
  OK
daisy: 
  crashed and auto rebooted again.
  Switched to other two fibers in cable, from green/blue to orange/brown.
marigold:
  running fine and dandy.

  Then at 7:10 pm CST (00:11 GMT) it went nuts, flooding the net 
  with traffic (removing its fiber from the hub quieted things down).
  Did not look at its console output.  A power cycle cured things.
ragwort:
  crashed and auto rebooted once or twice. When visited on tower,
  console was not responding, and status light was dull orange.
  Cable to Toshiba was loose, so no info was available.

  Rebooted ragwort with toshiba attached to console. After a while
  saw "med0: excessive collision" errors.

  Switched to other two fibers in cable, from orange/brown to green/blue.


Also, turned off the chem shelter air conditioner, in case the
problem could be power related.
The AC is on the same transformer as ragwort,daisy,marigold.







26: Status, Site base, Sat 25-Sep-1999 13:30:11 CDT, Daily Status
ISFF/CASES99 Daily status 25 Sept 1999

Weather: windy, cloudy in AM, clear midday, overcast in PM
Surface: dry

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   up      up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        intermittent crashes
marigold:       up
daisy:          up
cosmos:         up-fixed krypton crashes


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Awaiting Dave Miller (Univ. Conn.) to bring sensors.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.



25: Radiation, Site , Sat 25-Sep-1999 13:29:36 CDT, Pyg IC started!
The pyrgeometer intercomparison can be considered to have officially
started at 12:15 (1715 Z) today.  Yesterday, Tony mounted fans on
all housings.  Today, I just connected the radiometers which were
recalibrated (after their swim in oil) and received here last night.
(It was necessary to reposition both radiometers on their 8-hole base
plates to fit the housings.)

I also just leveled all of the Schwartz's (not knowing if this had
been done already) using the old digital level.  Obviously, I didn't
have a good reference surface, but I still think that I got them
to within 0.1 degree.  (They all sight to the horizon similarly.)

24: ADAMS, Site tower, Sat 25-Sep-1999 12:21:34 CDT, ADAM status
cosmos:
  Fixed its problems this morning.  Software bug in krypton calibration
  function couldn't handle empty lines in the cal file and would
  infinitely loop, hanging whatever host was running covar, cockpit, or
  prep.  Cosmos is now ingesting all its channels.
daisy: 
  rebooted itself this morning for no apparent reason.  
marigold:
  running fine and dandy.
ragwort:
  On Friday, 9/24, replaced the 40m fiber borrowed from ISS with
  a 4-fiber 300m cable.  The 40m cable is still tie-wrapped to the tower.
  This was done to fix problem losing fiber link (LMON light on Dyna FL/4
  port would go out and ragwort would reboot).

  ragwort died this morning.   The halt light on the Ironics card was red.
  The light on the Matrix was dim red I believe.  Power glitch?
  Rebooted, with the toshiba logging the console.


23: Radiation, Site , Fri 24-Sep-1999 10:59:27 MDT,
Replaced K&Z 970377 with K&Z 940187
 
Now have K&Z 940181     12.08 uV/Wm-2    82.78 Wm-2 / uV
         K&Z 940187      9.93 uV/Wm-2   100.70 Wm-2/ uV
22: Tower, Site tower, Fri 24-Sep-1999 09:06:49 MDT, 55m tower guy wire tensions

    55 meter tower guy wire tensions:


   September 13, 1999  (just after retensioning)

Guy level    Nominal     NW corner     SW corner     SE corner     NE corner
             Tension
              (lbs.)

   18'     425-475        475           475           575           450

   36'     325-375        260           160           300           350

   54'     225-275        200           250           250           215

   78'     575-625        690           510           675           625

   102'    525-575        600           525           700           600

   126'    625-675        650           675           625           625

   150'    575-625        650           510           525           700

   174'    725-775        750           725           775           650

   186'    675-725        590           675           725           675


   September 22, 1999  (weekly recheck)

                          NW            SW            SE            NE

   18'     425-475        450           450           425           370

   36'     325-375        265           275           310           280

   54'     225-275        180           190           250           200 

   78'     575-625        650           475           600           525

   102'    525-575        550           515           650           540

   126'    625-675        570           650           540           600

   150'    575-625        625           530           500           670

   174'    725-775        700           700           700           650

   186'    675-725        520           685           735           625


differences in guy tensions at the same level and at different times could
be caused by many different factors - wind loading, slipage of guy clamps,
slipage of anchors due to wind or moisture, etc.


 September 29, 1999 weekly recheck

                          NW            SW            SE            NE

   18'     425-475       375           400           390           290
   36'     325-375       260           250           300           300
   54'     225-275       200           180           230           210
   78'     575-625       600           440           500           500
   102'    525-575       510           540           525           615
   126'    625-675       570           600           520           550
   150'    575-625       580           500           495           625
   174'    725-775       725           680           670           600
   186'    675-725       550           600           630           625

 October 8, 1999 weekly recheck by Marcel V.
			NW		SW		SE		NE
18'	425-475		490		425		400		300
36'	325-375		250		250		300		450
54'	225-275		280		175		250		200
78'	575-625		575		475		540		460
102'	525-575		525		510		500		590
126'	625-675		550		610		525		550
150'	575-625		600		500		490		625
174'	725-775		700		675		700		600
186'	675-725		500		690		700		610


 October 18, 1999 weekly check by Steve S. (unit 2150)
			NW		SW		SE		NE
18'	425-475		410		400		425		290
36'	325-375		240		275		325		300
54'	225-275		200		175		240		200
78'	575-625		600		500		550		500
102'	525-575		550		575		550		575
126'	625-675		550		650		575		550
150'	575-625		575		525		500		650
174'	725-775		700		700		725		625
186'	675-725		500		700		675		600

 October 26, weekly check by Tony D and Sean  ? Wind from S (Unit 2150)
                        NW              SW              SE              NE
18'     425-475         340             380             400             280
36'     325-375         300             270             300             330
54'     225-275         200             180             260             230
78'     575-625         540             460             520             470
102'    525-575         480             520             510             560
126'    625-675         500             600             560             510
150'    575-625         550             520             500             610
174'    725-775         660             660             710             590
186'    625-725         510             710             660             600     






21: LOG, Site , Thu 23-Sep-1999 07:45:52 MDT, Boom Angles
Below is a table of each sensors' angle.
Angle is direction boom is pointing.

Angles were also shot on 10/23, 10/23

60 meter tower:

sensor		angle
--------------------------------------
ATIK (10m)	96 44' 35"
PROP (15m)	96 58' 10"
ATIK (20m)	96 44' 35"
PROP (25m)	96 44' 35"
CSAT (30m)	96 44' 35"
PROP (35m)	96 35' 10"
ATIK (40m)	96 44' 35"
PROP (45m)	96 44' 35"
CSAT (50m)	96 21' 55"
ATIK (55m)	96 04' 35"

---------------------------------------
5 meter tower:
---------------------------------------
CSAT (2.5m)
CSAT (5m)	96 27' 20"

---------------------------------------
PAM STATIONS
station sensor		angle
stn1	sonic		93 45 43	26 Oct		Cathy
stn1	prop		 0 48 05	26 Oct		Cathy
stn2	sonic	        95 05 23	28 Oct		Cathy
stn2	prop           359 37 40	28 Oct		Cathy
stn3	sonic		88 55 15	27 Oct	        Cathy
stn3	prop	       358 46 35	27 Oct		Cathy
stn4	sonic	        89 18 37	27 Oct	        Cathy
stn4	prop           356 63 00	27 Oct		Cathy
stn5	sonic           92 32 20	28 Oct 		Cathy
stn5	prop	       358 35 00	28 Oct		Cathy
stn6	sonic		88 06 20	26 Oct		Cathy
stn6	prop	       356 54 06	26 Oct		Cathy











20: Kryptons, Site , Tue 21-Sep-1999 07:16:55 MDT, krypton locations on 60 meter tower
Serial number 1101 was installed at 5 meter height.
The calibration coefficients are:
path -   1.011cm
Kw   -  -0.137
V0   -   2677

Serial number 1133 was installed at 55 meter heigth.
The calibration coefficients are:
path -   1.434
Kw   -  -0.153
V0   -   2886
19: PAM-EVE, Site , Mon 20-Sep-1999 08:12:54 MDT, PAM Sensor / Station ID Locator
PAM Sensor Locations at Setup and start of Ops
----------------------------------------------

Toucan Ports
   cua/0	stn1
   cua/1	stn2
   cua/2	stn3
   cua/3	stn4
   cua/4	stn5
   cua/5	stn6

Sensor		ST1	ST2	ST3	ST4	ST5	ST6
------		---	---	---	---	---	---
RMY - id	314	313	315	319	245	355
    - prop	62082	53664	48627	44341	43554	53686
    - pitch	307	307	304	309	309	309
    - BoomDir	002	0	358	357	0	357	

TRH		007	010	009	008	012	011

Sonic - id	NUW5	NUW7	NUW3	ATI3	NUW6	ATI2
    - Ht in cm	504	502	496	509	506	503
    - Azim.	See Below
    - Samples	10	10	10	10	10	10

Krypton		1390	1395	1391	1394	1392	1393
    - hasDeltaP	no	no	no	yes	yes	yes
Tsfc		3389-3	3577-3	3577-2	3577-6	3389-1

CR10		tbd	tbd	tbd	n/a	tbd	n/a
    - MRI Rain	13	3	32	n/a	n/a	n/a

Batt-PV		Conc.	Conc.		Dekka	Conc.	ABS
    - cont.	Morn.	Morn.	Morn.	Flex	Morn.	Morn.
    - box-id	3			New	5	1
							Elec.Fence

EVE - fmt	73	73	73	74	73	74

Notes:
(1) Sonic Boom Azimuths
	1 = 273:31:20
	2 = 271:57:57
	3 = 268:05:55
	4 = 270:01:10
	5 = 273:03:57
	6 = 268:26:23 check this, jwm
(2) All stations running w/20-hz sonic and krypton/delta-P.
(3) Delta-Pressure monitoring is declared in the krypton.dat file.
    There is no local storage on EVE of these data, only fastout.
    Download file from Rocky: c:\pam\cases99\krypton\no_dp\kryp13xx.dat
    for stations 1-3 which do not have the deltaP, 
    or c:\pam\cases99\krypton\with_dp\kryp13xx.dat for stations
    4-6 which do.  Then copy kryp13xx.dat to 'krypton.dat' on EVE.
(4) EVE Format 73 includes logger, Format 74 does not have logger.
    Both formats include gps clock reset every 3-hours, averages
    and covars with file output every 5-minutes to ramdisk metppf
    file.
(5) EVE prom 990926 is installed in all boxes.  Allows gps clock
    reset and 1/day 'cardcopy' of file output so that the
    traditional second file output process to /card is not needed.








18: BaseSystems, Site , Sun 19-Sep-1999 08:08:45 MDT, Changes in CONFIG files
I have made changes to prep, channel, archive, and
covar config files.
(1) 5 channels have been added to COSMOS to accommodate
the PYG intercomparison. COSMOS serial channels 211 thru
215 were used for these sensors. Gordon will need to come
up with a new cockpit display for COSMOS because there are
too many data coming in to show all the data. The designation
for these data are:
	channel		name (s0x: x is pam station number)
 	211		boom
	212		s01
	213		s02
	214		s03
	215		s05

(2) The 30 meter radiation boom was mounted at 50 meters.
So, the data for this sensor will go thru RAGWORT. The
designation for the sensor's data has been changed from
.30m to .50m. The data is going into channel 206.


 

17: LOG, Site , Sun 19-Sep-1999 07:48:19 MDT, STATUS - Sept 19
 The 5 lower booms now have sensors on, and data are
being collected. Hopefully today we will get another
3 or 4 up.
 Daisy seems to be acting better since the sensors
have been added. It was rebooted yesterday and stayed
up all night.
 Ragwort is still up adn down.
 Four PAM stations are set up with sensors except for
Tsurface. The prop/vanes are giving bad data. This may be
due to incorrect configuration of output units.

UPDATE:
3 more sensors added at 35m, 40m, and 45m. All TRHs in
DC power cables at daisy and marigold replaced with
shorter cables. All breakout boxes and 4 channel cables
installed at ragwort.
NOTE: daisy serial data down. This may be due to the new
power cable.



16: PropVanes, Site , Sun 19-Sep-1999 07:45:49 MDT, Tower Prop Pitchs
The table below provides the prop pitch information
for each propeller vane on the 60 meter tower

prop location	prop s/n	pitch
---------------------------------------------
15 meter	56182		304
25 meter	54606		304
35 meter	54614		304
45 meter	54636		304



15: LOG, Site , Sat 18-Sep-1999 09:45:58 MDT, STATUS - Sept 18
  All the sensor booms have been mounted except for
the 35 meter prop. Its location has a conflict with
the set of crossed tower guy wires.
(1) 5 of the TRH sensors have been mounted, 55 meter remains.
(2) 10 meter sonic mounted.
(3) 15 meter prop mounted.
(4) the pyg radiation intercomparison has been setup.
(5) 4 pam stations are installed, having problems with sonics
(6) Ragwort is still not working correctly
(7) The 50 meter radiation mounting bracket was tested on the
    5 meter boom track. Everything worked ok. We will try
    mounting the sensor for testing on a nice day.  


14: SoilSensors, Site , Wed 15-Sep-1999 12:20:17 MDT, Installed soil sensors
Between 10:30 and 12:30, Wed, 15 Sept, I installed a set of soil sensors:
     Soil Ht Flux Plate
     Soil Temperature Probe
     Soil Moisture Probe
at Station 1 / PAM 1, the Short grass site
   Station 2 / PAM 2, the Bluestem site
   Station 5 ? PAM 5, the Aromatic weed site

In the afternoon the sensors were wetted down with several gallons of water

Between 9:300 and 10:00, Fri, 17 Sept, the sensors were installed at
Sta 3 / PAM 3, the Green grass site. These sensors were also wetted down.

13: LOG, Site , Tue 14-Sep-1999 18:19:24 MDT, STATUS - Sept 14
(1) 4 booms installed
(2) cosmos on the air
(3) more PAM installation done
(4) started work on 5 meter boom
	will install at ~ 50 meters
(5) ragwort still having problems

weather: windy and cloudy today

12: SoilSensors, Site , Tue 14-Sep-1999 11:00:19 MDT, Soil sensor installation
Installation of soil sensors

Campbells and associated sensors to be installed at equivalently 
numbered PAM stations

           Ht Flux Plate      Soil T Probe        Soil Moisture Probe

Sta 1         87036             NCAR 001              001

Sta 2        H943149            NCAR 002              009

Sta 3        H983212            NCAR 003              006

Sta 4        H983213            NCAR 004              008

Sta 5        H983211            STP 97001             007

Sta 6         87034             STP 97003             005

Arrays for Sta 1, Sta 2, Sta 3 and Sta 5 installed at SE corner of Dark Horse


Heat Flux Plates at 5 cm depth
Soil T Probes at an angle from 1 cm to 4 cm
Soil Moisture probes at 2.5 cm depth.


11: LOG, Site , Mon 13-Sep-1999 07:57:31 MDT, STATUS: Sept 12
(1) 3 ADAMS up the tower
(2) tryed to get ragwort at 45 meters running
    some problem in communication
(3) toucan running
(4) more solar panels/batteries deployed at EVEs
(5) Tony located sites for rad stands
(6) Tony found turtle shell

weather: rain all morning, windy all day
         clear in the afternoon

10: LOG, Site , Sun 12-Sep-1999 09:32:36 MDT, STATUS - Sept 11
(1) most of the equipment was moved out to the tower
(2) solar panels and batteries moved to PAM stations
(3) removed plywood structure around tower
(4) started installation of Toucan
(5) started installation of first ADAM on tower

9: LOG, Site , Sat 11-Sep-1999 07:58:22 MDT, STATUS - Sept 10
At the end of the day the following items were done:
(1) base station located and powered up
(2) equipment unloaded from base and transport vehicle
(3) computer systems setup and running
(4) fiber optic line run from base to tower
(5) started solar panel deployment to PAM stations


8: Kryptons, Site , Fri 03-Sep-1999 13:03:55 MDT, How to Change PAM Krypton Calibration Files
How To Change PAM Station Configuration when swapping Kryptons
between stations:
--------------------------------------------------------------
The steps basically amount to updating the "krypton.dat" file on
the PAM station by using the Rocky computer.
It is optional whether or not to also update the records on the base.


0) Physically swap in the new sensor.

1) Update the PAM station "krypton.dat" file
   Use the rocky to interact with the station:
      EVE> entersys
      System # xmodem -yrt
      'On-Rocky' send appropriate file such as 'Kryp1393.dat'
      System EVE# copy kryp1393.dat krypton.dat
      System EVE# start
   And verify it works checking the id and data EVE sees.

   Important Note: the Rocky is setup with the appropriate files to
   send but you must select the proper version to use.  This is
   because the krypton files include a declaration for ingesting
   the 20-hz delta-pressure sensor for stations 4,5 and 6.
   For stations 4,5,6:
        use the krypton file with deltaP included, select
        c:\pam\cases99\krypton\with_dp\kryp13xx.dat
   For stations 1,2,3:
        use the krytpon file without deltaP included, select
        c:\pam\cases99\krypton\no_dp\kryp13xx.dat
   The krypton files for EVE include serial numbers 1389 thru 1395.


2) Update the Base Station Records (THIS IS OPTIONAL AND NOT REALLY NEEDED.
   IT IS PREDICATED ON AN EARLIER METHOD OF CREATING PAM STATION CONFIGURATIONS
   WHICH HAD THE KRYPTON DECLARATIONS IMBEDDED IN THE MAIN FORMAT ITSELF):
   Edit file:  $PAM/config/projects/CASES99
   Change dated entry for the appropriate station and add a new dated entry
   for the Krypton being swapped in something like this:
    # STA_ID LAT        LONG    ELEVATION  START_TIME           STOP_TIME        KRYPTON_SN LN_V0  1/PATHL_KW  BIAS
    1       37:40:00.00 -96:43:00.00  460  08/25/1999,00:00:00  10/05/1999,11:55:00  1389   8.518 -5.200857  0.0
    1       37:40:00.00 -96:43:00.00  460  10/05/1999,12:00:00  12/01/1999,00:00:00  1393   8.387 -5.346163  0.0
    #------------
   Note: The date time must be formatted with 2-digit month/day/hour/min/sec
         fields as shown
   Note: You can always tweak the file time values later.

   Create new station configuration files:
   cd $PAM/projects/CASES99/sta_cfgs
   evemake 73

   Note: this assumes you are using configuration format #73.  That may or
   may not change during the project.
   Note: evemake creates files called:  s01_0073.dat.990825
   for each station, format# and start date.
   These files can alternately be copied to the Rocky and the
   PAM station and renamed to 'config.dat', however the procedure
   outlined above should be more straightforward.





7: Kryptons, Site , Fri 03-Sep-1999 11:43:05 MDT, Krypton Calibrations for PAM / Config.dat File
Krypton calibrations dated from Campbell 8/25/99 added to: 
$PAM/config/projects/CASES99
For sensors 1389-1395.

PAM configuration files created assuming the following
station / sensor  associations for now:
1  1389
2  1390
3  1391
4  1392
5  1393
6  1394

See Also: How to change PAM Krypton Cal Files

6: ADAMS, Site none, Thu 02-Sep-1999 21:24:59 MDT, Datel sampling of 200 hz
Just noticed that the 200hz data from cosmos was actually coming in at
199hz.  The datel couldn't quite keep up, and was missing one
sample.  Cosmos is sampling 15 analog channels.  

The datel was computing averages of 8 samples when it is doing 17
channels or less.  If it does more than 17 channels, it averages
4 high-rate samples.  I lowered the cutoff from 17 to 14 channels.
Therefore it will be averaging 4 samples on cosmos, and 8 samples
on the other adams.



5: Kryptons, Site , Thu 02-Sep-1999 16:25:00 MDT, new kryptons tested
I just connected all of the new kryptons to cosmos:108 in serial number
(1389-1395) order.  This was done between 15:50 - 16:25 local.  All read
between 0.62 - 0.72 V, which was lower than what I expected, but still
usable.  The first one, 1389, I tried cleaning with de-ionized water, but
actually reduced the output from 0.72 to 0.68 V.  I cleaned it again and
got up to 0.74 V.  Thus, I didn't try cleaning any of the others.

4: PropVanes, Site , Mon 30-Aug-1999 10:46:38 MDT, PROP CALS
  Below are the propeller calbrations for the units used in
CASES99. The propellers on the PAM sensors have a higher pitch
then specified by RMYoung. As a test, one PAM propeller was tested
on an ASTER sensor.The computed pitch was 294. The correct value.
At this time, we are assuming the physical dimensions of the PAM
sensor has an effect on the flow resulting in the higher pitch values.
The raw data can be extracted from the computed data if we decide the
pitch values used are incorrect.
  In some cases, a propeller has already been assigned to a sensor.
This is noted in the table below.
PAM:
propeller	pitch (mm/rev)	sensor
53686		309		355
62082		307		314
44341		309		319
43554		309		245
53678		304		318
53664		307		313
48627		304		315
53665		306

ASTER
propeller	pitch (mm/rev)
56182		304
54614		304
54636		304
56159		306
53560		305
54631		304
56153		306
54606		304

3: BaseSystems, Site , Thu 29-Jul-1999 08:57:32 MDT, Sysadmin task list
toucan:
  headless, internal disk only: done 7/29
  solaris: core package with basic networking, ppp, ntp, yp
  latest patches: done 7/29

  priority paging done
  mount disks from aster: automount/NIS done


cocklebur:
  latest patches (2.7) done 7/29
  mount disks from aster: automount/NIS

clover:
  backup existing system to restore when clover returns
 
a replacement of clover for tony:


aster:
  NIS

manus: ppp system at ANL
  headless IPC with magma serial card
  to boot IPC from CDROM:  boot sd(0,6,2)
  install solaris: done 7/29
  install magma card from aster & driver

  install standalone ssh, tcpwrappers

framemaker?
Splus 5?








2: ADAMS, Site , Fri 18-Jun-1999 13:43:34 MDT, cosmos quirky
Cosmos (#940211101) started acting funny on 6/16.  It would seldom 
finish the reboot sequence successfully, or would die soon after booting.
Removing the datel did not help.  This was in backplane #4,
datel #440690.

Swapped in daisy (#9010) and all worked fine.

On 6/18 swapped cosmos back in and it seems to be working again...

It ran until 18:00 6/19, and died. The error message had scrolled of the
procom screen because of many "write to datel pipe" errors.

On Monday 6/21, it would hang during the startup sequence because it
couldn't talk to the Datel's DPR, failing to get the appropriate response
back after trying to start the Datel executive.





1: LOG, Site none, Fri 18-Jun-1999 13:42:13 MDT, Logbook created
initial log file /net/aster/projects/CASES99/logbook/tklog.log created by USER maclean